### v1

# Disentangling the galactic and intergalactic components in 313 observed Lyman-alpha line profiles between redshift 0 and 5

Abstract
Context. Lyman-Alpha (${\rm Ly}\alpha$) photons emitted in star-forming regions inside galaxies experience a complex radiative transfer process until they reach the observer. The ${\rm Ly}\alpha$ line profile that we measured on Earth is, thus, the convolution of the gas properties in the interstellar (ISM), circumgalactic (CGM), and intergalactic medium (IGM).
Aims. We make use of the open source package zELDA (redshift Estimator for Line profiles of Distant Lyman-Alpha emitters) to disentangle the galactic and IGM components of the ${\rm Ly}\alpha$ profiles to study both the evolution of the intrinsic galactic emission and the IGM transmission across cosmic time.
Methods. zELDA includes different artificial neural networks that reconstruct IGM attenuated ${\rm Ly}\alpha$ line profiles. These models are trained using mock ${\rm Ly}\alpha$ line profiles. A Monte Carlo radiative transfer code computes the galactic component for the so-called thin shell model. Moreover, the IGM component is included through the IGM transmission curves generated from the IllustrisTNG100 cosmological galaxy formation simulation.
Results. We recover their intrinsic galactic spectra by applying the zELDA to 313 ${\rm Ly}\alpha$ line profiles observed with HST/COS and MUSE. Sources at z ¡ 0.5 show weak IGM attenuation, while at z ¿ 3, ZELDA reveals significant IGM suppression of the blue peak in several sources. After separating the IGM effects, the stacked intrinsic galactic ${\rm Ly}\alpha$ profiles show a minimal evolution from z = 0 to 6. The mean IGM transmission for z ¡ 0.5 in HST/COS data exceeds 90%, while the MUSE data show an evolution from 0.85 at z = 3.0 to 0.55 at z = 5.0.
Key Words.:
Radiative transfer, Galaxies: intergalactic medium, Galaxies: ISM
1 Introduction
Figure 1: Illustration of the impact of the IGM attenuation in the galactic ${\rm Ly}\alpha$ line profile. Through all the panels, we show the intrinsic ${\rm Ly}\alpha$ line profile emerging from the galaxy in thin black and the IGM transmission curve in yellow. In the top panels, we show the galactic line profile obscured (purple). The central left panels show the galactic line profile after traveling a short distance in the IGM (blue). As the ${\rm Ly}\alpha$ travel further through the IGM the line is more and more absorbed in the blue side (top right central panel), until they reach the observer (red). The bottom panels show the zELDA reconstruction (purple) using as input the observed line profile (red), along with the line structure terminology used in this work.
The first spectral line of neutral hydrogen, Lyman-$\alpha$ (${\rm Ly}\alpha$), plays a pivotal role in extragalactic astronomy (Ouchi et al., 2020). It is not only used to detect and characterise the most distant galaxies (e.g. Zitrin et al., 2015; Witstok et al., 2024) but thanks to its resonant nature, it probes regimes otherwise inaccessible to observations, such as the dim outskirts of galaxies (Steidel et al., 2010; Wisotzki et al., 2016).
Specifically, the shape of the observed ${\rm Ly}\alpha$ line profile is the product of the complex radiative transfer processes that take place in the interstellar, circumgalactic, and intergalactic mediums (ISM, CGM, and IGM, respectively; for a review on the radiative transfer see Dijkstra, 2017). The high cross-section between neutral hydrogen and ${\rm Ly}\alpha$ causes these photons to be absorbed and reemitted, often thousands of times before leaving the galaxy. Each time a photon goes through one of these events, it changes in direction and frequency because of the velocity distribution of the hydrogen atoms. After escaping, the ${\rm Ly}\alpha$ line profile (frequency distribution) is greatly influenced by the gas geometry, density, and bulk velocity.
The fact that ${\rm Ly}\alpha$ observables probe gas from the ISM to the IGM holds tremendous potential. As such, ${\rm Ly}\alpha$ spectra are frequently used to probe the porosity of the ISM and as a proxy for ionizing photon escape (Izotov et al., 2016; Dijkstra et al., 2016; Verhamme et al., 2017, e.g.,). Furthermore, the existence of a dominant red peak in the ${\rm Ly}\alpha$ spectra (e.g. Erb et al., 2014; Trainor et al., 2015) is interpreted as a widespread presence of galactic outflows.
On the scale of the CGM, ${\rm Ly}\alpha$ reveals the gas properties through the shape of the surface brightness profiles on scales of tens of kpc for individual (e.g., Wisotzki et al., 2016) and even hundreds of kpc for stacked observations (Steidel et al., 2011; Wisotzki et al., 2018; Lujan Niemeyer et al., 2022). In addition, thanks to the advent of integral field spectrographs such as MUSE (Bacon et al., 2010) or KCWI (Martin et al., 2010), one can now also study of the emergent ${\rm Ly}\alpha$ spectrum changes as a function of projected position (e.g. Erb et al., 2018; Leclercq et al., 2017) revealing, for instance, the transition from galactic out- to intergalactic inflows (Erb et al., 2018; Li et al., 2022; Guo et al., 2024).
On even larger scales, the ${\rm Ly}\alpha$ photons emerging from the galaxy are shaped by the IGM, which contains neutral hydrogen in (mostly self-shielding) lumps. Thus, further radiative transfer processes influence the ${\rm Ly}\alpha$ line profile. However, because photons interacting with the IGM gas change direction and are scattered out of the line of sight – and the probability of scattering (back) into the line of sight is negligible – one can treat these large-scale effects as effective absorption. Importantly, in this process, photons are redshifted progressively by the Hubble flow as they travel through the IGM. This causes the IGM to attenuate higher frequencies of the galactic spectrum as photons advance (e.g. Laursen et al., 2011a).
We illustrate this in the top panels of Fig.1. The galactic ${\rm Ly}\alpha$ line profile is shown in black, and the IGM transmission is drawn in yellow. In the middle left panel (blue), we show a line profile after traveling a short distance in the IGM, which is slightly absorbed. Then, in the upper right middle panel (green), we show the exact line profile after continuing travel through the IGM, which shows more attenuation. Finally, in the top left panel (red), we show how the line reaches the observer.
The fact that the IGM also changes the ${\rm Ly}\alpha$ line shape and the observability of ${\rm Ly}\alpha$ emitting galaxies leads to the fact that ${\rm Ly}\alpha$ emission at $z\gtrsim 5$ is now regularly used as a powerful tool to constrain the ‘Epoch of Reionization’ (for a review, see, Dijkstra, 2014). While predominantly the (non-)visibility, as well as the observed strength of the ${\rm Ly}\alpha$ emission (through the change of the luminosity function, the equivalent width distribution, or through the clustering of LAEs), is used to constrain the (global) neutral fraction, it is also possible to use the observed spectral shape to constrain the neutral gas of the IGM. For instance, the most blueward observed flux can be used to constrain the size of the ionized region around a given galaxy (Mason & Gronke, 2020; Tang et al., 2024), in particular the visibility of an observed blue peak (Matthee et al., 2018; Gronke et al., 2020).
While this ability of ${\rm Ly}\alpha$ to probe neutral hydrogen on the ISM, CGM, and IGM scales provides the community with a potential treasure trove of information, this highly multiscale process naturally comes with many degeneracies. For instance, does the absence of a blue peak imply an ISM outflow or a neutral IGM? To unravel this degeneracy, we implemented the ability to reconstruct the transmission and the ‘galactic’ ${\rm Ly}\alpha$ spectra (shaped by the ISM and CGM) into the Python package zELDA  [^1] Github: https://github.com/sidgl/zELDA_II. Tutorials and documentation on zELDA are also available at https://zelda-ii.readthedocs.io/index.html. (redshift Estimator for Line profiles of Distant Lyman-Alpha emitters). To fit observed ${\rm Ly}\alpha$ line profiles, zELDA is now able to disentangle between the galactic and the IGM components of a ${\rm Ly}\alpha$ line profile through artificial neural network models trained with mock ${\rm Ly}\alpha$ spectra (see, Gurung-Lopez et al. (in prep.) for a detailed description and testing of the algorithm or Sect. 2 for a summary).
How the ISM, CGM, and IGM affect the observed ${\rm Ly}\alpha$ spectra and the ability to disentangle the ‘galactic’ from the intergalactic effects are visualized in Fig. 1. Here, we show in the top panels how the ${\rm Ly}\alpha$ line is affected by absorption in the IGM (as described above) and in the bottom panels how zELDA manages to reconstruct the ‘galactic’ line shape relatively well.
This work is structured as follows: in Sect. 2 we briefly introduce zELDA’s pipeline for recovering the galactic component from IGM attenuated ${\rm Ly}\alpha$ line profiles. We introduce the observational data set used in this work in Sect. 3. Next, we apply zELDA’s methodology to the observational data set in Sect. 4. Finally, we draw our conclusions in Sect. 5.
In this work, we draw the intergalactic transmission curves and ${\rm Ly}\alpha$ line profiles in the rest frame difference to the ${\rm Ly}\alpha$ wavelength ($\Delta\lambda_{0}$). Redshift accuracy is provided in the same units. Typically, this quantity is expressed in velocity units, instead, as $\Delta v=c\Delta\lambda_{0}/\lambda_{\rm Ly\alpha}\approx(247\,{\mathrm{km}\,\mathrm{s}^{-1}})\times\Delta\lambda/1$Å, where $c$ is the void speed of light and $\lambda_{\rm Ly\alpha}\approx 1215.67$Å.
Furthermore, we refer to the ${\rm Ly}\alpha$ IGM escape fraction of a source as the ratio between the flux emerging from the ISM/CGM and that observed after the IGM along the observer’s line of sight. This definition can be understood as the transmission of IGM around ${\rm Ly}\alpha$ and through the line of sight. In general, in the IGM, ${\rm Ly}\alpha$ photons are not destroyed by dust grains, although they are scattered out of the line of sight. This effectively causes a reduction of the observed ${\rm Ly}\alpha$ flux.
2 Reconstructing attenuated Lyman-$\alpha$ emission lines
In this section, we detail our methodology to reconstruct observed ${\rm Ly}\alpha$ line profiles attenuated by the intergalactic medium and, simultaneously, estimate their ${\rm Ly}\alpha$ IGM escape fraction. Our methodology is based on the Python package zELDA. Here, we briefly introduce zELDA’s pipeline, which is fully detailed in ZP24.
To fit the observed ${\rm Ly}\alpha$ line profiles, we use the open-source Python package zELDA (Gurung-López et al., 2019a, 2021, 2022). zELDA’s thin shell model is integrated by linear interpolation in a precomputed grid of line profiles provided by the full Monte Carlo radiative transfer code LyaRT (Orsi et al., 2012). The grid contains 3,132,000 nodes in the five dimensional space of bulk expansion velocity $V_{\rm exp}$ $\in[0,1000]kms^{-1}$, neutral hydrogen column density $N_{\rm H}$ $\in[10^{17},10^{21.5}]cm^{-2}$, dust optical depth $\tau_{a}$ $\in[0.0001,0.0]$, the intrinsic equivalent width $EW_{\rm in}$ $\in[0.1,1000]\AA{}$ and line width $W_{\rm in}$ $\in[0.01,6]\AA{}$ of the ${\rm Ly}\alpha$ emission before entering the thin shell. Further information on the grid and its accuracy can be found in Gurung-López et al. (2022).
The IGM transmission curves used to train the IGM+z and IGM-z models are described in detail in Byrohl & Gronke (2020); Byrohl et al. (2021). These are computed in the IllustrisTNG100 simulation (Naiman et al., 2018; Nelson et al., 2019; Marinacci et al., 2018; Pillepich et al., 2018; Springel et al., 2018). The ${\rm Ly}\alpha$ radiative transfer was computed using an updated version of the code ILTIS (Behrens et al., 2019; Byrohl & Gronke, 2020; Byrohl et al., 2021). An IGM transmission curve is obtained for every halo heavier than $5\times 10^{9}M_{\odot}$ in 1000 different lines of sight. This analysis was performed in six redshift snapshots at $z=$0.0, 1.0, 2.0, 3.0, 4.0 and 5.0.
zELDA includes three models based on artificial neural networks. For all of them, the input contains an observed ${\rm Ly}\alpha$ line profile. The models IGM+z and IGM-z take into account the IGM attenuation and provide an unattenuated best-fitting shell model. IGM+z is conceived as a model with the IGM redshift evolution of IllustrisTNG100. Meanwhile, IGM-z is built to be redshift independent. The third model, NoIGM, does not consider the IGM and fits a shell model directly to the observed line profile.
[…]
  • IGM+z: The training set includes mock line profiles attenuated by the IGM. The IGM transmission curve associated with a particular ${\rm Ly}\alpha$ line profile uses the actual redshift of the source. By construction, the IGM evolution of Byrohl & Gronke (2020); Byrohl et al. (2021) is directly imprinted on the training set. Thus, ${\rm Ly}\alpha$ line profiles at lower redshift are less attenuated than those at high redshift. Finally, IGM+z includes a proxy redshift in the input.
  • IGM-z: The training set also includes mock line profiles attenuated by the IGM. Although, in contrast to IGM+z, we use the re-calibrated  [^2] As shown in ZP24, Byrohl & Gronke (2020) IGM transmission curves were fine-tuned so that they match the mean evolution of the redshift of the IGM transmission by Faucher-Giguère et al. (2008). Thus, the IGM transmission curves used in IGM+z and IGM-z are slightly different. transmission curves. These are drawn randomly without considering the mock source redshift. Thus, there is no evolution of the IGM through redshift in this training set. There will be ${\rm Ly}\alpha$ basically unattenuated (typical of $z=0$) and greatly attenuated (typical of high $z$) at all redshifts. Finally, unlike IGM+z, IGM-z does not include the proxy redshift in the input.
  • NoIGM: The training set does not use IGM transmission curves. NoIGM includes a proxy redshift in the input, as in IGM+z.
The output of the three models is basically the same: the shell model parameters with their uncertainty, the redshift of the source, and the IGM escape fraction 2Å around the ${\rm Ly}\alpha$ wavelength, $f_{\rm esc}^{4\AA}$ (in the case of IGM+z and IGM-z).
The motivation for using IGM+z and IGM-z is to verify the validity of our results. While IGM+z is trained with the evolution of IGM transmission with redshift of IllustrisTNG100, IGM-z lacks it. Comparison between these two models can highlight possible biases in our methodology in redshift-dependent trends, such as the evolution of the IGM escape fraction. In principle, if analyzing IGM+z predictions, we find a redshift dependence trend missing from IGM-z data, then it could be induced by our training set rather than being real. Meanwhile, we show the results of NoIGM, a model without IGM attenuation for comparison.
Figure 2: Nine examples of zELDA’s prediction on line profiles displayed in the rest frame. The observed line profile is shown in dark grey with its $1\sigma$ uncertainty in light grey. zELDA’s reconstruction using the models IGM+z, IGM-z and NoIGM are displayed in blue, green and yellow, respectively. The redshift of the source is shown in the top right corner. zELDA’s estimations of $f_{\rm esc}^{4\AA}$ given by the IGM+z and IGM-z models are shown in blue and green, respectively. The central grey-shaded region shows the wavelength interval of $f_{\rm esc}^{4\AA}$. The estimated outflow parameters are indicated in the table with each subplot. $V_{\rm exp}$ is given in  km s-1, $N_{\rm H}$ in $cm^{-2}$ and $EW_{\rm in}$ and $W_{\rm in}$ in Å. All the fitted line profiles are also shown in Appendix C.
3 Description of observational data
The observational ${\rm Ly}\alpha$ line profiles used in this work are stored in the Lyman alpha Spectral Database (LASD, Runnholm et al., 2021) [^3] http://lasd.lyman-alpha.com, which currently contains 349 ${\rm Ly}\alpha$ emission lines between redshift 0 and 6.0. We remove from our analysis
  1. 1.
    ${\rm Ly}\alpha$ line profiles with a very steep continuum around ${\rm Ly}\alpha$,
  2. 2.
    spectra with low $S/N_{p}$   and
  3. 3.
    spectra that apparently contain the ${\rm Ly}\alpha$ emission line of several sources.
This leaves a total of 313 ${\rm Ly}\alpha$ line profiles from redshift 0 to 6.0. All the spectra are displayed in Appendix C.
The ${\rm Ly}\alpha$ line profiles from LASD were obtained mainly from two experiments: i) the Cosmic Origins Spectrograph (COS Green et al., 2012) on board the Hubble Space Telescope (HST) and ii) the MUSE-WIDE survey Urrutia et al. (2019); Herenz et al. (2017). ${\rm Ly}\alpha$ line profiles obtained via HST were observed in the General Observers (GO): GO 11522 and 12027 (PI: Green, Salzer et al., 2001; Wofford et al., 2013), GO11727 and 13017 (PI: Heckman, Heckman et al., 2011, 2015), GO 12269 (PI: Scarlata, Songaila et al., 2018), GO 12583 (PI: Hayes Hayes et al., 2014; Rivera-Thorsen et al., 2015), GO12928 (PI: Henry Henry et al., 2015), GO 13293 and 14080 (PI: Jaskot Jaskot & Oey, 2014; Jaskot et al., 2017), GO 14201 (PI: Malhotra Yang et al., 2017) and GO 13744 (PI: Thuan Izotov et al., 2016, 2018, 2020).
HST-COS and MUSE-WIDE exhibit very different spectral quality. On one side, the spectra obtained by HST have excellent spectral quality ($R\sim 16,000$). In particular, the ${\rm Ly}\alpha$ line profiles exhibit excellent wavelength sampling, as $\Delta\lambda_{\rm Pix}$ might take two particular values: 0.0598Å and 0.0735Å. This is given by the two medium resolution gratings G130M and G160M used to observe the sources. The spectral resolution is also high quality. $W_{\rm g}$ and ranges from 0.073Å to 0.10Å with median 0.085Å. The signal-to-noise ratio of the maximum of the line, $S/N_{p}$, spawns a wide range from $S/N_{p}$ =6.5 to $\sim 400$ with a median of $\sim$38 and only a $\sim 11\%$ of sample exhibit $S/N_{p}$ $<15.0$. On the other hand, considering MUSE data, the spectral quality is more limited ($R\sim 3000$), $\Delta\lambda_{\rm Pix}$ =1.25Å. Meanwhile, $W_{\rm g}$ changes from 1.25Å at $z=3$ to 4.0Å at $z=6.0$. We only consider ${\rm Ly}\alpha$ line profiles such as $S/N_{p}$5.0. The maximum $S/N_{p}$ is 35.0, and the median is 7.8.
In ZP24 several spectral quality configurations are explored. Most of the line profiles obtained by HST have a similar quality to the best-case scenario studied in the previous sections, and these ${\rm Ly}\alpha$ line profiles are presented in sections 4.2 and 4.3 in ZP24. Meanwhile, MUSE spectral quality is similar to the middle and worst cases in the same sections of ZP24. It is shown in ZP24 that for MUSE and HST spectral quality, zELDA can reconstruct the stacked ${\rm Ly}\alpha$ line profile, and in particular if there is evolution, as well as the $f_{\rm esc}^{4\AA}$ redshift evolution.
4 Results on observed Lyman-$\alpha$ line profiles
In this section, we apply zELDA’s ANN models to observed ${\rm Ly}\alpha$ line profiles. In Sect. 4.1 we display the predicted ${\rm Ly}\alpha$ line profiles and in Sect. 4.2 we compare the shell model parameters of the different ANN models. Later, in Sect. 4.3 we study zELDA’s accuracy to provide the source redshift in Sect. 4.5. We present our observed $\langle f_{\rm esc}^{4\AA}\rangle$ reconstruction in Sect. 4.6 and we compare it to the global ${\rm Ly}\alpha$ escape fraction in Sect. 4.7 and the intrinsic stacked ${\rm Ly}\alpha$ line profile in Sect. 4.8. We explore the fact that the IGM+z and IGM-z models predict that the ${\rm Ly}\alpha$ line profiles are IGM attenuated at $z\sim 0$ in Sect. 4.9. Finally, we discuss possible biases that might arise from fitting IGM attenuated line profiles with models lacking the IGM component in Sect. 4.10.
4.1 Individual reconstructed ${\rm Ly}\alpha$ line profiles
Figure 3: Comparison of the ANN output variables in the observed ${\rm Ly}\alpha$ line profiles predicted by different models. The out given by IGM+z is given in the X-axis and IGM-z in the Y-axis. HST ($z<0.55$) sources are shown in blue, while MUSE ($z>2.9$) sources are shown in red. The black dashed line marks the one-to-one relation. The Spearman coefficient for the full sample is shown in black (top), for HST sources in blue (middle), and for MUSE in red (bottom). The red and blue solid lines are the best-fitting first-degree polynomials for the high and low redshift samples. In the top row $V_{\rm exp}$, $N_{\rm H}$ and $\tau_{a}$ are shown from left to right. In the bottom row $EW_{\rm in}$, $W_{\rm in}$ and $f_{\rm esc}^{4\AA}$ are shown from left to right. The mean uncertainty for the shown properties is displayed in the bottom right corner.
In Fig. 2, we show nine examples of the performance of zELDA in observed ${\rm Ly}\alpha$ line profiles. The observed ${\rm Ly}\alpha$ line profile is shown in grey, while the predictions of IGM+z, IGM-z, and NoIGM are shown in blue, green, and yellow. The top row shows the HST ${\rm Ly}\alpha$ line profiles (sources A, B and C) while the middle and bottom rows show MUSE LAEs (sources from D to I ).
In general, IGM+z, IGM-z, and NoIGM reproduce the red peak of the observed ${\rm Ly}\alpha$ line profiles well. IGM+z and IGM-z predict a blue peak while differing from NoIGM. The NoIGM predicted ${\rm Ly}\alpha$ line profiles closely follow the observed line profile. This is the case for the HST and MUSE spectrum.  [^4] Note that, in mock ${\rm Ly}\alpha$ line profiles, a high $z$ observed line profile without a blue peak is not always reconstructed by IGM+z and IGM-z with a blue peak.
Figure 4: Same as for Fig. 3 but comparing IGM+z (X-axis) against NoIGM (Y-axis). NoIGM does not predict a value for $f_{\rm esc}^{4\AA}$, so we the comparison between NoIGM and IGM+z for $f_{\rm esc}^{4\AA}$ is not available.
Figure 5: Rest frame difference between zELDA’s ${\rm Ly}\alpha$ wavelength and the estimated ${\rm Ly}\alpha$ wavelength using other spectral feature different from ${\rm Ly}\alpha$ in HST sources. zELDA’s predictions are shown blue, green and yellow for IGM+z, IGM-z and NoIGM, respectively. The mean and standard deviation of the distributions are shown for each model.
At low redshift (top row of Fig. 2), the three ANN models (IGM+z, IGM-z, and NoIGM) predict nearly identical intrinsic line profiles for sources A and B, and the predicted profiles align closely with the observed spectra, particularly in both blue and red peaks. This consistency reflects the minimal IGM attenuation at these redshifts. In source C ($z=0.36$), all models predict similar red peak amplitudes, but the blue peak exhibits slight differences: IGM-z predicts the strongest blue peak, followed by IGM+z and then NoIGM. These differences correlate with variations in the predicted parameters, such as $V_{\rm exp}$, which decreases from 96.4 km s-1 (NoIGM) to 71.6 km s-1 (IGM+z) and 51.6 km s-1 (IGM-z). This trend indicates that the inclusion of IGM attenuation affects the reconstructed expansion velocity and suggests that IGM-z provides a more accurate fit by incorporating additional IGM constraints.
At higher redshift (middle and bottom rows), the influence of IGM absorption becomes more pronounced. For sources D, E, G, H, and I, IGM+z and IGM-z consistently predict a reconstructed blue peak, albeit attenuated or partially absorbed, while NoIGM either predicts no blue peak or one with significantly lower amplitude. For instance, in source I, the suppressed blue peak predicted by IGM+z and IGM-z ($f_{\rm esc}^{4\AA}$ $\sim 0.56$) matches the observations, whereas NoIGM fails to replicate this feature, highlighting the necessity of IGM modeling in accurately capturing line profiles at high redshift. Furthermore, the small difference in $f_{\rm esc}^{4\AA}$ values (within 5%) between IGM+z and IGM-z underscores the robustness of their predictions.
The observed and reconstructed ${\rm Ly}\alpha$ line profiles of all the sources are shown in Appendix C. Fig. 18 for sources with $0.200.0<z<$ 0.20, Fig. 19 for sources with $0.350.20<z<$ 0.35, Fig. 20 for sources with $3.30.35<z<$ 3.3 , Fig. 21 for sources with $3.73.3<z<$ 3.7 , Fig. 22 for sources with $4.23.7<z<$ 4.2 ,Fig. 23 for sources with $4.84.2<z<$ 4.8 , Fig. 24 for sources with $6.04.8<z<$ 6.0. In general, the red peak of the ${\rm Ly}\alpha$ predicted by the ANN models follows the observations closely. The blue peak predictions follow the observations at low redshift, while at high redshift, they diverge for IGM+z and IGM-z.
4.2 Comparison of Outflow Properties across Models
In Fig. 3, we compare the outflow properties and $f_{\rm esc}^{4\AA}$ predicted by the IGM+z (X-axis) and those predicted by IGM-z (Y-axis). HST sources (low redshift) are shown in blue, while MUSE sources (high redshift) are shown in red. We find that, in general, IGM+z and IGM-z predict the same outflow properties and $f_{\rm esc}^{4\AA}$. The typical Spearman correlation coefficient for the whole sample (shown in the top left corner in black) is above 0.8. For the low and high redshift samples, the output variables also correlate. We find that both models behave almost the same for $V_{\rm exp}$, $N_{\rm H}$, $\tau_{a}$, and $EW_{\rm in}$ at high and low redshifts. Meanwhile, at high redshift, the $W_{\rm in}$ and $f_{\rm esc}^{4\AA}$ values predicted by IGM+z and IGM-z are close to the one-to-one relation. Also, in the low redshift sample, for $W_{\rm in}$, although the Spearman correlation coefficient is high (0.82), the $W_{\rm in}$ values predicted by IGM-z are $\sim 0.25\AA$ above the IGM+z predictions. Finally, for $f_{\rm esc}^{4\AA}$ the low redshift Spearman correlation coefficient is also high (0.71), but IGM-z predicts slightly lower values than IGM+z.
In Fig. 4, we show a comparison between the output variables of IGM+z (X-axis) and those predicted by NoIGM (Y-axis). In general we find that there is a Spearman correlation coefficient higher than $\sim 0.7$ for the whole sample. The Spearman correlation coefficient is greater in the low redshift sample than in the high redshift sample. For example, for $V_{\rm exp}$, $\rho_{z<3}=0.72$ while $\rho_{z>3}=0.69$ and for $N_{\rm H}$ $\rho_{z<3}=0.83$ while $\rho_{z>3}=0.64$. The IGM+z model predicts higher values for $V_{\rm exp}$ at high redshift. Meanwhile, IGM+z predicts lower values of $W_{\rm in}$ than NoIGM at high redshift.
The results in Fig. 4 reveal systematic biases in shell model fits that do not account for IGM absorption (e.g. Gronke, 2017). These models overestimate expansion velocities by $\sim 0.30$dex and neutral hydrogen column densities by $\sim 0.25$dex. This occurs because the IGM primarily absorbs the blue side of the line – hence, when models exclude the IGM, they compensate by increasing both the galactic absorption’s blueshift ($V_{\rm exp}$) and the strength ($N_{\rm H}$).
Figure 6: Left: Distribution of the Kolmogórov-Smirnov (KS) estimator in the red tail of HST ${\rm Ly}\alpha$ line profiles for IGM+z, IGM-z, and NoIGM in thick blue, medium green, and narrow yellow, respectively. The width of the distribution is shown for each distribution, computed as the difference between the percentiles 84 and 16, and the median.
Figure 7: Comparison of the ANN output and LASD variables in the observed ${\rm Ly}\alpha$ line profiles predicted by the same models. $EW_{\rm obs}$ and $L_{\rm Ly\alpha}$ were obtained from LASD. $V_{\rm exp}$ is given in  km s-1, $N_{\rm H}$ in $cm^{-2}$, $EW_{\rm in}$, $EW_{\rm ob}$ and $W_{\rm in}$ in Å and $L_{\rm Ly\alpha}$ in $erg/s$. The blue dots show the values predicted by IGM+z while the green dots show that of IGM-z. The grey dots show the properties that include measurements by LASD. The histograms on the diagonal show the distribution of the parameters in the corresponding color for each model.
[…]
4.3 Accuracy of redshift estimates with zELDA
In this section, we study the redshift accuracy of zELDA. For this goal, we make use of the HST sources with a systemic redshift measured with a different feature than ${\rm Ly}\alpha$. We focus on HST sources as MUSE sources in LASD lack a systemic redshift below 0.5. This makes a total of 111 sources. To quantify the precision of the redshift, we analyze the distribution of the difference in rest frame wavelength between the predicted redshift by zELDA $z^{\rm zELDA}$ and the systemic redshift $z^{\rm Sys}$, computed as in Gurung-López et al. (2022), i.e.,
$\Delta\lambda^{\rm zELDA}_{0}=\lambda_{\rm Ly\alpha}\frac{z^{\rm zELDA}-z^{\rm Sys}}{1+z^{\rm Sys}},$ (1)
shown in Fig. 5. We find that zELDA provides a sub-angstrom accuracy using only the ${\rm Ly}\alpha$ line profile.
Our analysis reveals that both IGM+z and IGM-z outperform NoIGM in estimating the redshift of the sources, with IGM-z showing slightly lower dispersion compared to IGM+z. The comparison highlights the robustness of the IGM-inclusive models in reducing uncertainty and providing consistent predictions.
However, a small systematic bias is present across all models, with zELDA generally overestimating the ${\rm Ly}\alpha$ rest frame by approximately 0.25Å (equivalent to  61 km s-1). This overestimation aligns with findings from previous studies, such as Orlitová et al. (2018) and Li et al. (2021), and points to inherent limitations of the ‘thin shell model.’
Figure 9: $f_{\rm esc}^{4\AA}$ as a function of redshift for IGM+z (left) and IGM-z (middle). The right panel shows the histogram of the predicted $f_{\rm esc}^{4\AA}$ for IGM+z and IGM-z in blue an green, respectively.
4.4 Line profile fitting accuracy
To compare the models fairly, the accuracy of the recovered shape of the intrinsic ${\rm Ly}\alpha$ line profile in the observed spectrum, we focus on their red tails. In principle, galaxies at low redshift should be little absorbed by the IGM. In particular, most of the absorption should happen bluewards ${\rm Ly}\alpha$, leaving mostly untouched the red side of ${\rm Ly}\alpha$. Thus, the peak of the red peak and its tail extending redwards should be close or identical to that of the ${\rm Ly}\alpha$ line profile before traveling through the IGM. Therefore, comparing the red tail of the observed ${\rm Ly}\alpha$ line profiles to the predicted intrinsic spectrum in the same wavelength range should be a fair comparison. We define the red tail of our ${\rm Ly}\alpha$ line profiles as the wavelength range between the maximum of the red peak and ${\rm Ly}\alpha$ +5Å.
In Fig. 6, we show the distribution of the KS estimator for the IGM+z (blue), IGM-z (green), and NoIGM (yellow) in the red tails of the HST sources. We find that IGM-z and IGM+z outperform NoIGM. In particular, the median of the KS distribution for the models are 0.07 for IGM+z, 0.07 for IGM-z and 0.09 for NoIGM. Meanwhile, NoIGM exhibit a little bit more dispersion than IGM+z and IGM-z. We conclude that the IGM+z and IGM-z models predict the red tail of the observed ${\rm Ly}\alpha$ line profile with the same precision, being IGM-z only slightly better than IGM+z.
Figure 10: Mean $f_{\rm esc}^{4\AA}$ as a function of redshift (squares) for IGM+z (blue) and IGM-z (green). The solid lines indicate the fits, including all the sources. Meanwhile, the colored dashed line indicates the fit using only MUSE data. The grey dashed line shows the Faucher-Giguère et al. (2008) mean IGM transmission.
Table 1: Best-fitting parameters for the mean ${\rm Ly}\alpha$ IGM escape fraction ($\langle f_{\rm esc}^{4\AA}\rangle$) as a function of redshift, modeled with the modulated Fermi-Dirac distribution given by Equation 2.
$z$ range Experiments ANN model $a$ $b$ $c$
0.0-5.0 HST+MUSE IGM+$z$ 5.35$\pm$0.0143 0.8$\pm$0.0084 0.93$\pm$0.0002
0.0-5.0 HST+MUSE IGM-$z$ 5.5$\pm$0.0366 0.86$\pm$0.0196 0.89$\pm$0.0002
2.9-5.0 MUSE IGM+$z$ 5.18$\pm$0.187 0.66$\pm$0.084 1.0$\pm$0.033
2.9-5.0 MUSE IGM-$z$ 5.25$\pm$0.326 0.66$\pm$0.133 1.0$\pm$0.057
4.5 Correlations between model and observed properties
In this section, we study the existence of a correlation between the thin shell outflow properties predicted by zELDA, redshift and the ${\rm Ly}\alpha$ luminosity $L_{\rm Ly\alpha}$ and rest frame observed ${\rm Ly}\alpha$ equivalent width $EW_{\rm ob}$. The latter two quantities are obtained directly from LASD.
In Fig. 7, we show the predicted properties against all the other properties. $V_{\rm exp}$, $N_{\rm H}$, $\tau_{a}$, $EW_{\rm in}$, $W_{\rm in}$, $z$, $f_{\rm esc}^{4\AA}$, $EW_{\rm ob}$ and $L_{\rm Ly\alpha}$ from left to right and from top to bottom. The top right triangle ( green and grey) shows the properties predicted by IGM-z against themselves. The bottom left triangle (blue and grey) shows those of IGM+z. The diagonal shows the distribution of the parameters.
Meanwhile, the Spearman correlation coefficient for the entire sample is shown in Fig. 8 with the $1\sigma$ uncertainty. The top left corner shows the Spearman correlation coefficient for IGM-z, $\rho_{-z}$(green). Meanwhile, the bottom left corner shows the Spearman correlation coefficient for IGM+z, $\rho_{+z}$(blue). To estimate the uncertainty proceed as in Gurung-López et al. (2022), we draw 1000 times a new set of outflow parameters from the posterior of the ANN models (see the methodology in ZP24). We compute the Spearman correlation coefficient for each set of parameters. The $1\sigma$ uncertainty is calculated by the percentiles 16 and 84 from the distribution of Spearman coefficients. We consider that two properties correlate or anti-correlate when the absolute value of the Spearman correlation coefficient is greater than 0.45. Furthermore, the Spearman correlation coefficients for the MUSE and HST separately are presented in Appendix B.
We find that, in general, the Spearman coefficient provided by the IGM+z and IGM-z models are similar. Typically, both models exhibit or lack the same correlations. Also, Fig. 17 shows the Spearman correlation coefficient for the low redshift (HST, left) and high redshift samples (MUSE right). Considering the complete ${\rm Ly}\alpha$ line profile sample, we find that only a handful of properties correlate. Overall, some of these correlations are also present in MUSE and HST samples by themselves, but others are missing.
We find the following relations:
[…]
4.6 IGM escape fraction evolution with redshift
Figure 11: Redshift evolution of different ${\rm Ly}\alpha$ escape fractions. The filled colored squares show the $\langle f_{\rm esc}^{4\AA}\rangle$ found by the IGM+z and IGM-z models up to $z=5.0$, while those empty show the results at $z>5$. The colored solid lines show the best fit to $\langle f_{\rm esc}^{4\AA}\rangle$ as presented previously. The dashed red line shows the global ${\rm Ly}\alpha$ escape fraction estimation by Hayes et al. (2011). The grey symbols show different studies from the literature measuring the global ${\rm Ly}\alpha$ escape fraction and their uncertainty. The colored dashed line shows the ISM escape fraction estimated from the combination of Hayes et al. (2011) and our results.
In this section, we explore in detail the $f_{\rm esc}^{4\AA}$ redshift dependence predicted by the models including an IGM component (igmz and IGM-z).
As shown in ZP24 (section 4.3 and appendix C), the accuracy of the IGM escape fraction $f_{\rm esc}^{4\AA}$ depends on the spectral quality of the ${\rm Ly}\alpha$ line profiles. In this section, we restrain our analysis to ${\rm Ly}\alpha$ line profiles with $S/N_{p}$8.0 and $W_{\rm g}$3.0Å. This results in a sample composed of 93 HST and 113 MUSE sources, totaling 206 ${\rm Ly}\alpha$ line profiles.
In Fig. 9, we show the $f_{\rm esc}^{4\AA}$ values for the 206 sources given by IGM+z (left) and IGM-z (middle), as well as their distributions (right). We find that $f_{\rm esc}^{4\AA}$ decreases with increasing redshift in both ANN models. In particular, $f_{\rm esc}^{4\AA}$ in HST sources ($z<0.55$) is concentrated between 0.8 and 1.0 for both IGM+z and IGM-z. Considering MUSE sources at $z>3.0$, $f_{\rm esc}^{4\AA}$ is concentrated between 0.85 and 0.5 for both IGM+z and IGM-z with dispersion $\sim 0.2$.
This intuitively makes sense, of course, as the higher neutral hydrogen content and increased IGM density at greater redshifts lead to stronger absorption of ${\rm Ly}\alpha$ photons, thereby reducing the IGM escape fraction.
In Fig. 10, we show the mean ${\rm Ly}\alpha$ IGM escape fraction, $\langle f_{\rm esc}^{4\AA}\rangle$ for the models IGM-z (blue squares) and IGM+z (green squares). The uncertainties were computed as the uncertainty of the mean. The bins have been shifted $\pm$0.01 in redshift for displaying proposes. We find that the $\langle f_{\rm esc}^{4\AA}\rangle$ predicted by IGM+z and IGM-z diverge at $z<0.5$ in HST data, while they match at $z>3.0$ in MUSE data. In HST data, the IGM+z and the IGM-z models predict no evolution in $\langle f_{\rm esc}^{4\AA}\rangle$ with redshift. IGM+z predicts a constant value of $\sim$0.91, while IGM-z predicts $\sim$0.89. Meanwhile, at $z>3.0$, both models predict that $\langle f_{\rm esc}^{4\AA}\rangle$ changes from 0.9 at $z=$2.9 to 0.6 at $z=5.0$.
In Fig. 10 we also show the best fitting of the predicted $\langle f_{\rm esc}^{4\AA}\rangle$ to a modulated Fermi-Dirac distribution (zELDA is able to recover the $f_{\rm esc}^{4\AA}$-$z$ evolution as discussed in ZP24), i.e.,
$\langle f_{\rm esc}^{4\AA}\rangle=\displaystyle\frac{c}{e^{b(z-a)}+1}.$ (2)
The best fitting parameters can be found in Tab.1. We find that fitting the whole redshift range (HST+MUSE) for IGM+z gives a best fitting $c=0.93$, while for IGM-z $c=0.89$. . However, in the $5.02.9<z<$ 5.0 range, the parametric curves are closer than than 0.05. Interestingly, when fitting only MUSE data ($5.02.9<z<$ 5.0), the best fitting curve for both ANN models are very similar. In fact, for both models $c=1.0$. Also, $a$ and $b$ are $1\sigma$ compatible. This causes a difference at $z<2.5$ in the best-fitting curves when considering the complete and only MUSE samples
For comparison, we also show $\langle T_{\rm IGM}\rangle$, the mean IGM transmission (Faucher-Giguère et al., 2008) in grey. The relationship between $\langle T_{\rm IGM}\rangle$ and $f_{\rm esc}^{4\AA}$ is a priory unknown: $\langle T_{\rm IGM}\rangle$ reflects the mean IGM transmission. In contrast, $f_{\rm esc}^{4\AA}$ measures the escape of ${\rm Ly}\alpha$ from LAEs on intermediate IGM scales and the ${\rm Ly}\alpha$ line profiles emerging from the ISM and CGM. For instance, if red peak dominated ${\rm Ly}\alpha$ line profiles enter into the IGM, the ${\rm Ly}\alpha$ IGM escape fraction $f_{\rm esc}^{4\AA}$ will be high compared to blue peak dominated ${\rm Ly}\alpha$ line profiles. $\langle f_{\rm esc}^{4\AA}\rangle$ predicted by IGM-z and IGM+z fall 0.1 below $\langle T_{\rm IGM}\rangle$. The difference between $f_{\rm esc}^{4\AA}$ and $\langle T_{\rm IGM}\rangle$ can be due to the stronger absorption around the ${\rm Ly}\alpha$ line center (Laursen et al., 2011b; Gurung-López et al., 2022). Even if the IGM transmission is $\sim 1$ a few Angstroms from ${\rm Ly}\alpha$, the stronger absorption at $-2\AA{}<\Delta\lambda<0.5\AA{}$ can affect the blue peak of $z=0$ lines. Yet, at high redshift ($z>1.5$), both, the parametric best fitting curves for both ANN models and the predicted $\langle f_{\rm esc}^{4\AA}\rangle$ are above $\langle T_{\rm IGM}\rangle$. While implying a non-trivial $\langle f_{\rm esc}^{4\AA}\rangle$-$\langle T_{\rm IGM}\rangle$ relationship evolution, this trend could also indicate a selection effect. In this work, we measure $f_{\rm esc}^{4\AA}$ in MUSE galaxies with observed strong ${\rm Ly}\alpha$ emission rather than the whole underlying galaxy population. Hence, it is likely that measuring $f_{\rm esc}^{4\AA}$ in ${\rm Ly}\alpha$ flux-limited samples might result in $\langle f_{\rm esc}^{4\AA}\rangle$ values above those of the whole star-forming galaxy population.
4.7 Comparison between the global and IGM escape fractions
The global ${\rm Ly}\alpha$ escape fraction, $f_{\rm esc}^{\rm Lya}$, is defined as the ratio between the total amount of ${\rm Ly}\alpha$ photons produced inside a galaxy and those reaching the observer. Thus, $f_{\rm esc}^{\rm Lya}$ encapsulates information about the escape of ${\rm Ly}\alpha$ from the ISM, CGM, and IGM. The redshift evolution of $f_{\rm esc}^{\rm Lya}$ has been explored in the literature (Blanc et al., 2011; Ouchi et al., 2008; Cassata et al., 2011; Sobral et al., 2018; de La Vieuville et al., 2019; Thai et al., 2023; Goovaerts et al., 2024). In particular, Hayes et al. (2011) calculated the $f_{\rm esc}^{\rm Lya}$ redshift evolution comparing the ${\rm H}\alpha$ flux, mostly unaffected by radiative transfer processes, to the measured ${\rm Ly}\alpha$ flux in the same sources.
In Fig. 11, we show the $f_{\rm esc}^{\rm Lya}$ redshift evolution obtained by Hayes et al. (2011) (dashed red line) along with the individual data points from the other studies in grey as described in the legend. Other studies, such as Goovaerts et al. (2024) (based on Thai et al. (2023)), also obtained relations between the redshift and $f_{\rm esc}^{\rm Lya}$ , finding compatible results with Hayes et al. (2011). Overall, it is found in the literature that $f_{\rm esc}^{\rm Lya}$ increases with redshift, being earlier galaxies less attenuated in ${\rm Ly}\alpha$. Normally, this is attributed to the dust in the ISM of galaxies, which is less abundant at high redshift Hayes et al. (2011); Gurung-López et al. (2019b); Goovaerts et al. (2024). In particular, the different experiments find that $f_{\rm esc}^{\rm Lya}<5\%$ at $z<3.5$. Also, at $65<z<$ 6, Ouchi et al. (2008); Cassata et al. (2011); de La Vieuville et al. (2019) found a $f_{\rm esc}^{\rm Lya}$ bellow 0.2 with relatively low uncertainty of $\sim 0.1$. Meanwhile, with lager uncertainties, Sobral et al. (2018) found a $f_{\rm esc}^{\rm Lya}\sim 0.3$ at $64<z<$ 6 and Goovaerts et al. (2024) found $f_{\rm esc}^{\rm Lya}\sim 0.8$ at $z=6$. Given the uncertainties, all these studies are compatible with each other.
Figure 12: Comparison between stacked ${\rm Ly}\alpha$ line profiles. The observed stacked ${\rm Ly}\alpha$ line profile is shown in grey, while zELDA’s prediction using IGM+z, IGM-z, and NoIGM are shown in blue, green, and yellow. The stacked ${\rm Ly}\alpha$ line profile is computed in 6 redshift bins. Each panel show the redshift bin indicated in their panel.
In this work, we have inferred the IGM escape fraction,$f_{\rm esc}^{4\AA}$, from the shape of the ${\rm Ly}\alpha$ line profiles. The global ${\rm Ly}\alpha$ escape fraction should always be lower than the ISM, CG,M and IGM escape fractions, as $f_{\rm esc}^{\rm Lya}$ is the product of these. Our direct measurements of $\langle f_{\rm esc}^{4\AA}\rangle$ (blue and green squares for IGM+z and IGM-z) are always above the measured $f_{\rm esc}^{\rm Lya}$ (grey dots), which is reassuring.
Interestingly, we find that the IGM escape fraction decreases with redshift. This is expected from a theoretical point of view given the following two facts. First, it is found by Hayes et al. (2011) and this work that the stacked galactic ${\rm Ly}\alpha$ line profile does not evolve significantly from redshift 0 to 6. Second, the IGM becomes denser at higher redshifts and absorbs more ${\rm Ly}\alpha$ photons (Faucher-Giguère et al., 2008; Laursen et al., 2011b; Gurung-López et al., 2020; Byrohl & Gronke, 2020). Moreover, we have included zELDA’s prediction on observed spectrum at $z>5$ as colored empty squares. $f_{\rm esc}^{4\AA}$ values for $z>5$ sources should be taken cautiously as the IGM transmission curves only reach $z=5.0$. We find that zELDA predicts that $\langle f_{\rm esc}^{4\AA}\rangle$ remains constant from $z=5$ to 6. This can happen since the blue peak of the ${\rm Ly}\alpha$ line is always absorbed, but the red peak is relatively little attenuated. The flattening in $\langle f_{\rm esc}^{4\AA}\rangle$ is also found for $z>4.0$ in the training set of IGM+z (see ZP24), but in the observed data IGM+z only predicts a stable $\langle f_{\rm esc}^{4\AA}\rangle$ from redshift $z=5$ onward. Given this, extrapolating the best fit of the $\langle f_{\rm esc}^{4\AA}\rangle$ (colored solid lines) to higher redshifts than $z=5$ is probably uncertain.
Moreover, we have estimated the ISM+CGM ${\rm Ly}\alpha$ escape fraction redshift evolution from the best fit of Hayes et al. (2011) global escape fraction and our IGM escape fraction. The mean ISM escape fraction, $f_{\rm esc}^{\rm ISM}$, is computed as $\langle f_{\rm esc}^{\rm Lya}\rangle/$$\langle f_{\rm esc}^{4\AA}\rangle$ and is displayed as the blue and green dotted line when using $\langle f_{\rm esc}^{4\AA}\rangle$ from IGM+z and IGM-z, respectively. The global ${\rm Ly}\alpha$ escape fraction is dominated by the ISM up to redshift $z=5$ when the ISM mean escape fraction matches $\langle f_{\rm esc}^{4\AA}\rangle$. Then, $f_{\rm esc}^{\rm ISM}$ grows until it reaches unity at $z=6$ (extrapolating the IGM+z and IGM-z best fits). If we consider that $\langle f_{\rm esc}^{4\AA}\rangle$ keeps constant from redshift $z=5$ onward, then the $f_{\rm esc}^{\rm ISM}$ would reach unity at $z=7$.
At the same time, our best fit for $\langle f_{\rm esc}^{4\AA}\rangle$ and for $f_{\rm esc}^{\rm Lya}$ by Hayes et al. (2011) cross at redshift 6 (or $z=7$ if the $\langle f_{\rm esc}^{4\AA}\rangle$ is kept constant from $z=5$). This is puzzling as the global ${\rm Ly}\alpha$ escape fraction should always be lower than that of the IGM component. Considering the case at high redshift in which the $f_{\rm esc}^{\rm ISM}=1$ due to the low dust abundance, only the IGM would obscure the ${\rm Ly}\alpha$ line. In this scenario, the global escape fraction should match that of the IGM.
We find some tension when extrapolating our results on the IGM escape fraction and those on the global escape fraction in the literature. On the one hand, Goovaerts et al. (2024)’s measurement at redshift $z=6$ suggests that the global ${\rm Ly}\alpha$ escape fraction could be compatible with unity already. At the same time, extrapolating Hayes et al. (2011), $f_{\rm esc}^{\rm Lya}$ would reach unity at $z\sim 11$. On the other hand, we find that the IGM escape fraction decreases with redshift up to $z=5.0$. Additionally, somewhere between redshift 5 and 11, the epoch of reionization takes place, meaning a higher abundance of neutral hydrogen in the universe, particularly in the IGM. This will cause a much lower $\langle f_{\rm esc}^{4\AA}\rangle$ if the typical ${\rm Ly}\alpha$ line profile remains the same, from $z=0$ to $z=6$.
This tension could be alleviated if the typical ${\rm Ly}\alpha$ line profile at $z>6$ is much more dominated by a red peak than those observed at lower redshifts. If the ${\rm Ly}\alpha$ radiation is redshifted enough within galaxies, then the IGM would have less impact. Thus, $\langle f_{\rm esc}^{4\AA}\rangle$ could increase at higher redshifts.
Another scenario is that neither extrapolating our estimation for the IGM escape fraction nor those like that of Hayes et al. (2011) is a valid option. Until now, the global ${\rm Ly}\alpha$ escape fraction has only been measured up to $z=6$. Measurements at higher redshifts can potentially further insights.
4.8 Stacked Lyman-alpha line profile fitting
In this section, we analyze the redshift dependence of the observed ${\rm Ly}\alpha$ stacked line profile shape (a similar analysis was done by Hayes et al., 2021a, 2023; we focus here on the reconstruction of the pre-IGM spectra using zELDA). We limit our analysis to ${\rm Ly}\alpha$ line profiles with $S/N_{p}$9.0. This ensures that more than $\sim$85of the ${\rm Ly}\alpha$ line profiles of HST are recovered with KS¡0.1. Meanwhile, as $W_{\rm in}$ changes from 1.25Å to 4Å in MUSE data, $\sim$85and $\sim$65of the line profiles should be recovered with KS¡0.1, respectively (see appendix D in ZP24). $S/N_{p}$9.0 was chosen as a compromise between reconstruction quality and number counts. We verify that similar results are obtained across different signal-to-noise thresholds ($S/N_{p}$6.0, 7.0, 8.0, 10.0), with higher thresholds producing noisier stacked line profiles, while the overall trends remain consistent.
We compute the stacked ${\rm Ly}\alpha$ line profile from a sample of observed ${\rm Ly}\alpha$ line profiles as follows. First, we apply zELDA’s ANN model (IGM+z, IGM-z or NoIGM) to the line profiles to obtain the reconstructed ${\rm Ly}\alpha$ line profiles and the redshift of the source. All the line profiles are shifted with respect to their rest frame in a fixed wavelength array. The individual ${\rm Ly}\alpha$ line profiles are normalized so that their maximum reaches unity. Then, the stacked line profile is computed as the median flux in each wavelength bin.
In Fig. 12, we present the ${\rm Ly}\alpha$ stacked line profiles across six redshift bins: $0.50.0<z<$ 0.5 (94 sources), $3.53.0<z<$ 3.5 (19 sources), and $4.03.5<z<$ 4.0 (17 sources) in the top row (left to right), and $4.54.0<z<$ 4.5 (10 sources), $5.04.5<z<$ 5.0 (13 sources), and $6.05.0<z<$ 6.0 (5 sources) in the bottom row (left to right). To ensure a fair comparison, the ${\rm Ly}\alpha$ profiles predicted by the ANN models are computed at the same resolution ($W_{\rm g}$) as the observations.
In general, we find that the observed stack spectrum (shown in grey in Fig. 12) changes with redshift. At $0.50.0<z<$ 0.5 we find a faint but detected continuum. Meanwhile, at higher redshift, no continuum is detected. At $0.50.0<z<$ 0.5, the ${\rm Ly}\alpha$ stacked line profile exhibits a clear double peak line profile, with a prominent red peak that extends $\sim 2\AA{}$ redder than ${\rm Ly}\alpha$. The blue peak is weaker and also extends $\sim 2\AA{}$ bluewards ${\rm Ly}\alpha$. The same structure of the double peak is also apparent at $3.53.0<z<$ 3.5. Thus, the ${\rm Ly}\alpha$ stacked line profiles for $0.50.0<z<$ 0.5 and $3.53.0<z<$ 3.5 exhibit only minor differences. At $4.03.5<z<$ 4.0, the blue peak becomes less pronounced and eventually disappears for $z>4.0$, while the red peak remains consistently present across all redshift bins. This is consistent with the findings by Hayes et al. (2021b).
Again, the observed stacked ${\rm Ly}\alpha$ line profile at redshift $0.50.0<z<$ 0.5 and that at $3.53.0<z<$ 3.5 exhibit a very similar blue peak. This is consistent with our measurements of the $\langle f_{\rm esc}^{4\AA}\rangle$ and the redshift invariance of the galactic stack line profile. We find that $\langle f_{\rm esc}^{4\AA}\rangle$ changes little from $z\sim 0.5$ to $z\sim 3.0$. Therefore, if the galactic stacked line profile is the same at $z\sim 0.5$ to $z\sim 3.0$, then the observed stack should be very similar at these redshifts. Reassuringly, this is what we find.
Regarding the ${\rm Ly}\alpha$ stacked line profiles predicted by the ANN models, we find that all three (IGM+z, IGM-z, and NoIGM) generally reproduce the observed profile redwards of the ${\rm Ly}\alpha$ wavelength. However, notable differences arise on the blue side. As expected, the NoIGM model closely follows the observed profile, showing a weak but noticeable blue peak at $z<4.0$ that nearly disappears at $z>4.0$.
Figure 13: Six examples of zELDA’s prediction on observed line profiles displayed in rest frame. The observed line profile is displayed on dark grey with its $1\sigma$ uncertainty in light grey. zELDA’s galactic line profiles of IGM+z, IGM-z and NoIGM are displayed in blue, green and yellow, respectively. The redshift of the source and the outflow properties for each models are shown in the corresponding color. $V_{\rm exp}$ is given in  km s-1, $N_{\rm H}$ in $cm^{-2}$ and $EW_{\rm in}$ and $W_{\rm in}$ in Å.The central grey shaded region shows the wavelength interval of $f_{\rm esc}^{4\AA}$. The red arrows show mark potential IGM features.
In comparison, IGM+z and IGM-z behave similarly. Overall, the ${\rm Ly}\alpha$ stacked line profile predicted by both ANN models remains unchanged through redshift. Both ANN models predict that the red peak becomes slightly narrower with increasing $z$, as found in the observations. However, both IGM+z and IGM-z predict that the amplitude of the blue peak is relatively constant and that it becomes slightly narrower at higher redshift.
4.9 IGM absorption in ${\rm Ly}\alpha$ line profiles at $z<0.5$
In Section 2 and 4, we have discussed the IGM+z and IGM-z models. These two models have very different training sets. On the one hand, IGM+z makes use of the IGM transmission curves from Byrohl & Gronke (2020) without the recalibration [^5] We refer here to the recalibration to the IGM transmission curves from Byrohl & Gronke (2020) so that they match the mean IGM redshift evolution by Faucher-Giguère et al. (2008). See ZP24. matching the redshift of the source.
On the other hand, IGM-z uses the IGM transmission curves after recalibration to match the mean IGM transmission in Faucher-Giguère et al. (2008), and the IGM redshift is randomized. These differences cause the $f_{\rm esc}^{4\AA}$ distribution to evolve differently with redshift in the training sets (see section 3 in ZP24). Thus, the $f_{\rm esc}^{4\AA}$ predictions are different between these models. For example, the IGM-z model gives an unbiased $f_{\rm esc}^{4\AA}$ measurement from $z=0$ to $z=5.0$ for mocks with different $\langle f_{\rm esc}^{4\AA}\rangle$ redshift evolutions with $S/N_{p}$ =15.0 and $W_{\rm in}$ =0.1Å. Meanwhile, for the same spectral quality, IGM-z overpredicts $\langle f_{\rm esc}^{4\AA}\rangle$ by a 5% when $\langle f_{\rm esc}^{4\AA}\rangle$ $\sim 0.8$ at $z<1.0$, while at $z>1.0$ $\langle f_{\rm esc}^{4\AA}\rangle$ is recovered correctly (see section 4.3 in ZP24).
Despite the differences between IGM+z and IGM-z, both models predict that the ${\rm Ly}\alpha$ line profiles at $z<0.5$ are attenuated by the IGM. IGM+z predicts a mean attenuation of $\sim 0.92$. Also, IGM-z predicts $\langle f_{\rm esc}^{4\AA}\rangle$ $\sim 0.89$. Both models predict individual values spawning from $\sim$0.8 to 1.0. If these lines are IGM attenuated, given the high signal-to-noise and resolution of HST ${\rm Ly}\alpha$ line profiles, a few individual absorbed systems should be visible.
In Fig. 13, we show six examples of ${\rm Ly}\alpha$ line profiles that could be showing IGM features at $z<0.5$. Example A is a double peak ${\rm Ly}\alpha$ line profile that has a very good signal to noise in individual pixels. At $\Delta\lambda_{0}\sim-3\AA{}$, in the tail of the blue peak, there is an apparent absorption. In contrast, IGM+z, IGM-z, and NoIGM predict that the tail of the blue peak decreases smoothly. This kind of feature is not present in the shell model. Thus, this might be an IGM absorption. Moreover, in example A, the red tail is quite well fitted by the three models. However, the red peak observed spectrum falls faster than the models. This could indicate that the IGM is absorbing the core of the line. This is also visible in examples C and D.
Also, in example B we find a plateau of flux $\sim 0$ at the core of the line (from $\Delta\lambda_{0}=\sim-1.5\AA{}$ to 0.3Å) [^6] Note that zELDA has problems reproducing these wide absorption droughts at line center since we restrict ourselves to $T=10^{4}\,$ gas. Higher (effective) temperatures lead to a wider absorption and thus can alleviate the need for central IGM absorption This is not the case for absorption features off line-center also marked in Fig. 13. . In comparison, the continuum of this source is detected around $0.05$. Additionally, this line shows a wide red peak that extends up to $\Delta\lambda_{0}\sim 4\AA{}$. Symmetrically, the blue peak seems to extend down to $\Delta\lambda_{0}\sim-4\AA{}$. As in case A, example B shows a decrease in flux in the blue peak at $\Delta\lambda_{0}\sim-3\AA{}$. Neither of these two features is found in the shell models and might be compatible with IGM absorption.
Furthermore, in cases C, D and F, we find that the blue peak exhibits small drops in flux. However, considering the signal-to-noise of individual pixels, these drops are significant. The three models IGM+z, IGM-z, and NoIGM match well the red peak. The models and observations match the ascension of the blue peak ($\Delta\lambda_{0}$ between -0.5Å and 0.0Å) well – especially in cases C and D. Also, the predicted model is slightly above the observations, while they mat again at the tail of the blue peak ($\Delta\lambda_{0}$ between -2.0Å and -1.8Å).
Finally, example E is again a double peak ${\rm Ly}\alpha$ line profile with very good signal-to-noise in individual pixels. We find that IGM+z and NoIGM models predict almost the same intrinsic line profile, while IGM+z predicts a line with the same red peak but with a more substantial blue peak. Focusing on the blue peak, we find that in the observed ${\rm Ly}\alpha$ line profile, there is more flux in the tail of the blue peak than in the IGM+z and NoIGM models. However, the blue tail is very well fitted by the IGM-z model. An absorbing system in the IGM could also cause this.
In summary, across examples A–F, potential signatures of IGM absorption are consistently observed in the blue peak, including small flux drops and deviations from the smooth profiles predicted by the models. Additionally, features such as steeper declines in the red peaks and wider absorption plateaus at line center further suggest the impact of IGM on ${\rm Ly}\alpha$ line profiles. These findings highlight the IGM’s role in shaping ${\rm Ly}\alpha$ profiles, particularly in regions where the ‘shell model’ struggles to fully replicate the observed spectra [^7] Note that other models of intragalactic ${\rm Ly}\alpha$ radiative transfer will have similar problems to carve out the steep absorption features observed as radiative transfer and the associated frequency redistribution generally smooths spectra..
Figure 14: Distribution of the difference between the input shell model properties and those recovered for the mock with HST-like quality. In the top panels $V_{\rm exp}$, $N_{\rm H}$ and $\tau_{a}$ are shown from left to right. In the bottom panels, $W_{\rm in}$, $EW_{\rm in}$, and $f_{\rm esc}^{4\AA}$ are shown from left to right. The prediction using IGM+z, IGM-z, and NoIGM are shown in blue, green, and yellow, respectively. The dashed horizontal lines mark the distribution’s mean matching the used model’s color. The black dashed line is set at 0 for reference. In top right corner a table summing up the mean ($\mu$) and standard deviation ($\sigma$) is displayed for the three models.
Figure 15: Same as Fig. 14, but for the mocks with the spectral quality of MUSE.
4.10 Possible biases when fitting an IGM attenuated line with a model without IGM
The shell model has been widely used to successfully reproduce observed ${\rm Ly}\alpha$ line shapes, providing valuable insights into the physical properties of ${\rm Ly}\alpha$-emitting systems. However, previous work has often neglected the impact of IGM absorption, which can systematically alter the observed line profiles and bias parameter estimates. In this section, we investigate the possible biases induced by fitting IGM-attenuated ${\rm Ly}\alpha$ line profiles with models that do not account for IGM effects.
In this section, we study the possible biases induced by fitting an IGM attenuated line with a model without IGM. For this goal, we use NoIGM as the model without IGM absorption. Meanwhile, we also compare the results for the explored models with IGM, IGM+z, and IGM-z.
The bias in the shell model parameters in a given ${\rm Ly}\alpha$ line profile population might depend on the distribution of actual shell model parameters intrinsic to the sample. To make our analysis realistic, we build mock ${\rm Ly}\alpha$ line profile samples from the shell model parameters measured in the observed ${\rm Ly}\alpha$ line profiles in Sect. 4. In particular, we produce a single line profile with the same shell model parameters as each of the observed line profiles with matching spectral quality. Then, we convolve the spectrum with an IGM transmission curve at the redshift of the source and measure the $f_{\rm esc}^{4\AA}$. If the measured $f_{\rm esc}^{4\AA}$ is within 1% of that original measure, we keep the IGM transmission curve. Otherwise, we draw another one until the criteria are met. This ensures that the $f_{\rm esc}^{4\AA}$ distribution in the mock is very similar to that obtained from the observed ${\rm Ly}\alpha$ line profile. As a result of this procedure, the sample size, redshift distribution, and true shell model parameters in our mock match those of the observed sample.
In Fig. 14, we show the distribution of the difference between the measured shell model parameters and the true ones for the mock line profiles of HST. Meanwhile, in Fig. 15, we show those with MUSE spectral quality. In the top right corner, we display the mean and standard deviation of the distributions. Ideally, if the methodology and spectral quality were perfect, the output should match exactly the input, thus all the differences should be null, and the mean and standard deviation should be zero.
Focusing on the HST-like sample (Fig. 14), the IGM+z and IGM-z models exhibit biases lower than 0.06 dex, in general. Meanwhile, the NoIGM model is unbiased at the level of IGM+z and IGM-z in $W_{\rm in}$ and $EW_{\rm in}$. In contrast, the mean of the distributions of NoIGM is biased towards lower values of $V_{\rm exp}$ (0.1 dex) and higher values of $N_{\rm H}$ (0.2 dex) and $\tau_{a}$ (0.3 dex).
Regarding the MUSE-like sample (Fig. 15), the IGM+z and IGM-z models are generally less biased than NoIGM. In particular, IGM+z and IGM-z exhibit a bias smaller than 0.1 dex in $V_{\rm exp}$, $\tau_{a}$, $EW_{\rm in}$, and $f_{\rm esc}^{4\AA}$. We find that IGM+z and IGM-z are biased 0.15 dex in $W_{\rm in}$. Moreover, we find that NoIGM is more biased than IGM+z and IGM-z, particularly in $V_{\rm exp}$ and $W_{\rm in}$. Interestingly, we find that NoIGM (0.15 dex) is slightly less biased than IGM+z and IGM-z ($\sim$0.23 dex) when estimating $N_{\rm H}$.
Comparing the HST-like and MUSE-like samples, we find that the NoIGM model is generally more biased than IGM+z and IGM-z. Additionally, we find that the standard deviation of the distribution of differences changes with the spectral quality. In general, we find that the width of the distribution is large in the MUSE-like sample. This direct effect of the lower spectral quality makes the line profile reconstruction more difficult.
5 Summary and conclusions
The observed ${\rm Ly}\alpha$ line profile is the result of the combination of the radiative transfer processes taking place in the interstellar, circumgalactic, and intergalactic mediums. In this work, we have disentangled the IGM from the ISM/CGM components in 313 observed ${\rm Ly}\alpha$ line profiles by using the python package zELDA.
The latest version of zELDA includes different artificial neural network models to recover the galactic component from an IGM attenuated ${\rm Ly}\alpha$ line profile. These models are trained using mock ${\rm Ly}\alpha$ lines computed using grid of precomputed ’shell model’ (Gurung-López et al., 2022) with a full Monte Carlo radiative transfer treatment (LyaRT, Orsi et al., 2012) and the ${\rm Ly}\alpha$ IGM transmission published by Byrohl & Gronke (2020) based on IllustrisTNG100 simulation (Nelson et al., 2019).
We have applied zELDA to 313 observed ${\rm Ly}\alpha$ line profiles, extracted from LASD (Runnholm et al., 2021). Of them, 111 come from HST/COS observations, and 202 from MUSE observations. zELDA’s results on observed ${\rm Ly}\alpha$ line profile are:
  • In general, we find that ${\rm Ly}\alpha$ line profiles are little attenuated in HST data (z¡0.5). Meanwhile, in MUSE data ($z>3.0$), both, IGM+z and IGM-z, predict that the blue side of ${\rm Ly}\alpha$ is attenuated. The ISM ${\rm Ly}\alpha$ line profiles predicted by our models frequently exhibit a stronger blue peak than that observed. This trend becomes prominent at higher redshift. This suggests that ${\rm Ly}\alpha$ line profile reconstruction becomes more important at higher redshifts. Therefore, galaxy kinematic studies using the ${\rm Ly}\alpha$ line profile at high redshift might benefit from using the IGM reconstructed line profile rather than the observed line profile directly.
  • We quantify the IGM escape fraction, $f_{\rm esc}^{4\AA}$, for the HST and MUSE samples. We find that the mean $f_{\rm esc}^{4\AA}$ at $z<0.5$ is different depending on the ANN model. While IGM+z predicts $\langle f_{\rm esc}^{4\AA}\rangle$ =0.92, IGM-z predicts $\langle f_{\rm esc}^{4\AA}\rangle$ =0.89. As discussed, IGM+z tends to be biased towards 1.0 in this redshift range, so the measurement provided by IGM-z should be more reliable. Moreover, both ANN models agree well when determining $\langle f_{\rm esc}^{4\AA}\rangle$ in MUSE data. IGM+z and IGM-z predict that $\langle f_{\rm esc}^{4\AA}\rangle$ evolve from 0.87 at $z=3.0$ to 0.55 at $z=5.0$. This measurement can be helpful to obtain a measured ’IGM-free’ ${\rm Ly}\alpha$ luminosity function. Interestingly, our models predict that $z<0.5$ ${\rm Ly}\alpha$ line profiles can exhibit IGM features such as those of the ${\rm Ly}\alpha$ forest. We have shown six potential examples of this.
  • We have compared the mean IGM ${\rm Ly}\alpha$ escape fraction predicted by zELDA and the global ${\rm Ly}\alpha$ escape fraction. We find that $f_{\rm esc}^{4\AA}$ is higher than the global ${\rm Ly}\alpha$ escape fraction in the range where both quantities are measured. Extrapolating, we find that at $z=6$, the IGM escape fraction could match the global escape fraction. This would imply that the ISM escape fraction would be unity at $z>6$. Thus, at higher redshifts the global ${\rm Ly}\alpha$ escape fraction should match that of the IGM.
  • Next, we have analyzed the ${\rm Ly}\alpha$ stacked line profile. In the observed stacked line profile, a red peak that becomes slightly thinner with redshift is accompanied by a weaker blue peak that faints progressively at higher redshift. This is in good agreement with previous studies (Hayes et al., 2023). Interestingly, the IGM+z and IGM-z models predict that the stacked ${\rm Ly}\alpha$ line profile before entering the IGM does not evolve strongly with redshift. In the galactic spectrum, the blue peak amplitude is almost invariant, and the line profile becomes slightly thinner at higher redshift. This would imply that the mechanism that shapes the galactic ${\rm Ly}\alpha$ line profile is invariant with redshift, at least up to $z\sim 5$. Thus, the redshift evolution in the observed ${\rm Ly}\alpha$ line profiles can be attributed to the radiative transfer in the IGM.
  • We examined the biases that arise when fitting an IGM-attenuated ${\rm Ly}\alpha$ line profile using a model that does not account for the IGM component. Some shell model parameters, like the intrinsic equivalent width, $EW_{\rm in}$, can be accurately retrieved. However, parameters such as the neutral hydrogen column density, $N_{\rm H}$, and the bulk velocity, $V_{\rm exp}$, may be biased depending on the spectral quality. This highlights the importance of taking into account the IGM absorption in the ${\rm Ly}\alpha$ line profiles at high redshift.
In times to come, zELDA could facilitate remarkable scientific explorations. For instance, with a substantial sample of LAEs and spectra of sufficient quality, zELDA may enable the 3D mapping of the IGM ${\rm Ly}\alpha$ escape fraction. This mapping could permit the analysis of cross-correlations between $f_{\rm esc}^{4\AA}$ and LAEs, potentially shedding light on whether the large-scale characteristics of the IGM are linked to ${\rm Ly}\alpha$ observability (Zheng et al., 2011).
Acknowledgements.
The authors acknowledge the financial support from the MICIU with funding from the European Union NextGenerationEU and Generalitat Valenciana in the call Programa de Planes Complementarios de I+D+i (PRTR 2022) Project (VAL-JPAS), reference ASFAE/2022/025. This work is part of the research Project PID2023-149420NB-I00 funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU. This work is also supported by the project of excellence PROMETEO CIPROM/2023/21 of the Conselleria de Educación, Universidades y Empleo (Generalitat Valenciana). MG thanks the Max Planck Society for support through the Max Planck Research Group. DS acknowledges the support by the Tsinghua Shui Mu Scholarship, funding of the National Key R&D Program of China (grant no. 2023YFA1605600), the science research grants from the China Manned Space Project with no. CMS-CSST2021-A05, and the Tsinghua University Initiative Scientific Research Program (no. 20223080023). This research made use of matplotlib, a Python library for publication quality graphics (Hunter, 2007), NumPy (Harris et al., 2020) and SciPy (Virtanen et al., 2020).
Appendix A $f_{\rm esc}^{4\AA}$ inference by an alternative model
Figure 16: Distributions of $f_{\rm esc}^{4\AA}$ predicted values for HST (left) and MUSE ${\rm Ly}\alpha$ line profiles (right). The distribution for REC+z, IGM+z, and IGM-z are shown in purple, blue, and green, respectively. The vertical dashed colored lines indicate the distribution median for the corresponding model.
[…]
We find the same trends in the performance of IGM+z, IGM-z, and REC+z in the observed spectrum as the mock spectrum. The distribution of the predicted $f_{\rm esc}^{4\AA}$ values of the observed ${\rm Ly}\alpha$ line profiles presented in Sect. 4 for REC+z (purple), IGM+z (blue,) and IGM-z (green) are shown in Fig. 16. The left panel shows HST sources at $z<0.55$, while the right panel shows MUSE sources at $z>2.9$. For MUSE sources, the three models show very similar $f_{\rm esc}^{4\AA}$ distributions with $\langle f_{\rm esc}^{4\AA}\rangle$ $\sim 0.7$ (horizontal dashed lines). Not only distribution in $6.02.9<z<$ 6.0 is the same, but also the redshift evolution, although it is not explicitly shown. Meanwhile, considering HST sources, the IGM+z and IGM-z exhibit very similar $f_{\rm esc}^{4\AA}$ distribution, with $\langle f_{\rm esc}^{4\AA}\rangle$ $\sim 0.89$ and $0.91$ respectively. Meanwhile, REC+z predicts higher $f_{\rm esc}^{4\AA}$ values, as the $f_{\rm esc}^{4\AA}$ distribution peaks at $\sim 1$ with small dispersion. Actually, the minimum $f_{\rm esc}^{4\AA}$ predicted by REC+z is $\sim 0.9$. Consequently, the $\langle f_{\rm esc}^{4\AA}\rangle$ for REC+z is $\sim 0.98$.
The training set in an artificial neural network determines its output fidelity. In particular, biases and trends inherent to the training set are imprinted on the output. We find that REC+z gives a $\langle f_{\rm esc}^{4\AA}\rangle$ of $\sim 1$ at $z<0.5$ in contrast to IGM+z and IGM-z that gives $\langle f_{\rm esc}^{4\AA}\rangle$ $\sim 0.9$. Meanwhile, the three models give very similar predictions at high redshift. Therefore, the low $z$ discrepancy comes from the training set of REC+z. The low $f_{\rm esc}^{4\AA}$ variance at $z<1$ in REC+z causes its predictions to be biased towards $f_{\rm esc}^{4\AA}$ =1 at low $z$, rather than being a real prediction.
Appendix B Property correlation in MUSE and HST
Figure 17: Same as Fig. 8 but only for HST sources in the left panel and MUSE sources in the right panel. The uncertainty is not shown for simplicity but is similar to the previous figure’s.
In Fig.17 we show the show the spearman correlation coefficients for the HST (left) and MUSE (right) samples. We find that in general, all the correlations found in the complete sample (see Fig.8) are also present in the individual samples, such as the correlation between $EW_{\rm in}$ and $EW_{\rm ob}$.
Other correlations existing in the combined sample are only present in one of them. For example, the anti correlation between $f_{\rm esc}^{4\AA}$ and redshift is present in the MUSE and combined sample. However, no trend is found within the HST sources.
Appendix C all fitted lines
In this appendix, we show all the fitted lines for both MUSE and HST data. Line profiles are sorted by redshift. Fig. 18 shows sources with $0.200.0<z<$ 0.20, Fig. 19 for sources with $0.350.20<z<$ 0.35, Fig. 20 for sources with $3.30.35<z<$ 3.3 , Fig. 21 for sources with $3.73.3<z<$ 3.7 , Fig. 22 for sources with $4.23.7<z<$ 4.2 ,Fig. 23 for sources with $4.84.2<z<$ 4.8 and Fig. 24 shows sources with $6.04.8<z<$ 6.0. The observed ${\rm Ly}\alpha$ line profile is shown in grey, while the prediction of IGM-z, IGM+z, and NoIGM are shown in blue, green, and yellow, respectively. Furthermore, $f_{\rm esc}^{4\AA}$ is shown in the top left corner in the same color as the model used, alongside the source redshift. The grey shaded band marks the wavelength region where $f_{\rm esc}^{4\AA}$ is estimated.
In Fig. 25, we show the ${\rm Ly}\alpha$ line profiles from LASD that were not considered for the analysis in this work. They removed these line profiles because i) they looked noisy or ii) they exhibited a very steep continuum not modeled in zELDA, or ii) the fits of IGM-z, IGM+z, and NoIGM looked unsuccessfully. From the original sample of 349 spectra, 35 (10%) were excluded.
In general, at low redshift (Fig. 18 and 19) IGM+z, IGM-z and NoIGM predict similar intrinsic ${\rm Ly}\alpha$ line profiles. In sources 7, 40 and 111 IGM+z and IGM-z predict almost the same ${\rm Ly}\alpha$ line profile and NoIGM predicts a different one. In sources 1, 14, 30, 38, 66, 82 and 97 IGM+z and IGM-z actually predict a different intrinsic ${\rm Ly}\alpha$ line profile. Also, in general, the $f_{\rm esc}^{4\AA}$ predicted by IGM+z and IGM-z are close to unity, being IGM-z $f_{\rm esc}^{4\AA}$ predictions slightly lower than those of IGM+z, in general.
Meanwhile, at higher redshift (Fig. 21, 22, 23, and 24), the intrinsic ${\rm Ly}\alpha$ line profile predicted by IGM+z and IGM-z match, while they diverge from that of NoIGM. Although in some cases, the three models match relatively well, such in sources 116, 119, 142, 162, 186, 220, and 299. Note that this scenario is more frequent at lower redshift, closer to $z=3.0$. Meanwhile, at higher redshifts, closer to $z=6.0$, there is more difference between the prediction and IGM+z/IGM-z and NoIGM. Also, in $\sim$80of the high redshift sources, the IGM+z and IGM-z predict the presence of a blue peak in the ${\rm Ly}\alpha$ intrinsic spectrum.
Figure 18: zELDA’s prediction on observed line profiles displayed in restframe. The observed line profile is shown in grey. zELDA’s reconstruction using the models IGM+z, IGM-z, and NoIGM are displayed in blue, green, and yellow, respectively. The redshift of the source is shown in the top left corner. zELDA’s estimations of $f_{\rm esc}^{4\AA}$ given by the IGM+z and IGM-z models are shown in blue and green, respectively. The grey-shaded region shows the wavelength interval of $f_{\rm esc}^{4\AA}$. The red number indicate the index in LASD.
Figure 19: Same as Fig. 18, but for other observed ${\rm Ly}\alpha$ line profiles between $z=0.2$ and $z=0.35$.
Figure 20: Same as Fig. 18, but for other observed ${\rm Ly}\alpha$ line profiles between $z=0.35$ and $z=3.3$.
Figure 21: Same as Fig. 18, but for other observed ${\rm Ly}\alpha$ line profiles between $z=3.3$ and $z=3.7$.
Figure 22: Same as Fig. 18, but for other observed ${\rm Ly}\alpha$ line profiles between $z=3.7$ and $z=4.20$.
Figure 23: Same as Fig. 18, but for other observed ${\rm Ly}\alpha$ line profiles between $z=4.3$ and $z=4.8$.
Figure 24: Same as Fig. 18, but for other observed ${\rm Ly}\alpha$ line profiles between $z=4.8$ and $z=6.0$.
Figure 25: Same as Fig. 18, but for other observed ${\rm Ly}\alpha$ line profiles for which exhibit low $S/N_{p}$ (e.g., 0 , 1 ), a steep continuum not modeled in zELDA (e.g., 0 , 1 ), or with for than two components (e.g., 3).