### v2

# Disentangling the galactic and intergalactic components in 313 observed Lyman-alpha line profiles between redshift 0 and 5

Abstract
[…]
1 Introduction
Figure 1: Illustration of the impact of the IGM attenuation in the galactic ${\rm Ly}\alpha$ line profile. Through all the panels, we show the intrinsic ${\rm Ly}\alpha$ line profile emerging from the galaxy in thin black and the IGM transmission curve in yellow. In the top panels we show the galactic line profile obscured intrinsically (top left, purple), the same profile after traveling a short distance in the IGM (blue, center left) as well as further through the IGM (green, top right), and its shape when it reaches the observer (red). The bottom panel show the zELDA reconstruction (purple) using as input the observed line profile (red), along with the line structure terminology used in this work.
The first spectral line of neutral hydrogen, Lyman-$\alpha$ (${\rm Ly}\alpha$), plays a pivotal role in extragalactic astronomy (Ouchi et al., 2020). It is not only used to detect and characterise the most distant galaxies (e.g. Zitrin et al., 2015; Witstok et al., 2024) but thanks to its resonant nature, it probes regimes otherwise inaccessible to observations, such as the dim outskirts of galaxies (Steidel et al., 2010; Wisotzki et al., 2016).
The emergent ${\rm Ly}\alpha$ line profile results from complex radiative-transfer processes in the interstellar, circumgalactic, and intergalactic media (ISM, CGM, and IGM; see Dijkstra 2017 for a review). Because of the large scattering cross-section of neutral hydrogen, ${\rm Ly}\alpha$ photons are absorbed and reemitted many times before escaping the galaxy. Each scattering event changes both the direction and the frequency of the photon according to the velocity distribution of the gas. Consequently, the observed ${\rm Ly}\alpha$ profile is strongly shaped by the geometry, density, and bulk motions of the gas.
The fact that ${\rm Ly}\alpha$ observables probe gas from the ISM to the IGM holds tremendous potential. As such, ${\rm Ly}\alpha$ spectra are frequently used to investigate the porosity of the ISM and as a proxy for ionizing photon escape (e.g., Izotov et al., 2016; Dijkstra et al., 2016; Verhamme et al., 2017). Furthermore, the existence of a dominant red peak in the ${\rm Ly}\alpha$ spectra (e.g. Erb et al., 2014; Trainor et al., 2015) is interpreted as a widespread presence of galactic outflows.
On the scale of the CGM, ${\rm Ly}\alpha$ reveals the gas properties through the shape of the surface brightness profiles on scales of tens of kpc for individual (e.g., Wisotzki et al., 2016) and even hundreds of kpc for stacked observations (Steidel et al., 2011; Wisotzki et al., 2018; Lujan Niemeyer et al., 2022). In addition, thanks to the advent of integral field spectrographs such as MUSE (Bacon et al., 2010) or KCWI (Martin et al., 2010), one can now also study the emergent ${\rm Ly}\alpha$ spectrum changes as a function of projected position (e.g. Erb et al., 2018; Leclercq et al., 2017), revealing, for instance, the transition from galactic out- to intergalactic inflows (Erb et al., 2018; Li et al., 2022; Guo et al., 2024).
At even larger scales, the ${\rm Ly}\alpha$ photons emerging from the galaxy are shaped by the IGM, which contains neutral hydrogen in (mostly self-shielding) lumps. Thus, further radiative transfer processes influence the ${\rm Ly}\alpha$ line profile. However, because photons interacting with the IGM gas change direction and are scattered out of the line of sight – and the probability of scattering (back) into the line of sight is negligible – one can treat these large-scale effects as effective absorption. Importantly, in this process, photons are redshifted progressively by the Hubble flow as they travel through the IGM. This redshifting causes the IGM to attenuate higher frequencies of the galactic spectrum as photons advance (e.g. Laursen et al., 2011a).
We illustrate this in the top panels of Fig.1. The galactic ${\rm Ly}\alpha$ line profile is shown in black, and the IGM transmission is drawn in yellow. In the middle left panel (blue), we show a line profile after traveling a short distance in the IGM, which is slightly absorbed. Then, in the upper right middle panel (green), we show the exact line profile after continuing travel through the IGM, which shows more attenuation. Finally, in the top left panel (red), we show how the line reaches the observer.
The fact that the IGM also changes the ${\rm Ly}\alpha$ line shape and the observability of ${\rm Ly}\alpha$ emitting galaxies leads to the fact that ${\rm Ly}\alpha$ emission at $z\gtrsim 5$ is now regularly used as a powerful tool to constrain the ‘Epoch of Reionization’ (for a review, see, Dijkstra, 2014). While predominantly the (non-)visibility, as well as the observed strength of the ${\rm Ly}\alpha$ emission (through the change of the luminosity function, the equivalent width distribution, or through the clustering of LAEs), is used to constrain the (global) neutral fraction, it is also possible to use the observed spectral shape to constrain the neutral gas of the IGM. For instance, the most blueward observed flux can be used to constrain the size of the ionized region around a given galaxy (Mason & Gronke, 2020; Tang et al., 2024), in particular the visibility of an observed blue peak (Matthee et al., 2018; Gronke et al., 2020).
While this ability of ${\rm Ly}\alpha$ to probe neutral hydrogen on the ISM, CGM, and IGM scales provides the community with a potential treasure trove of information, this highly multiscale process naturally comes with many degeneracies. For instance, does the absence of a blue peak imply an ISM outflow or a neutral IGM? To unravel this degeneracy, we implemented the ability to reconstruct the transmission and the ‘galactic’ ${\rm Ly}\alpha$ spectra (shaped by the ISM and CGM) into the Python package zELDA  [^1] Github: https://github.com/sidgl/zELDA_II. Tutorials and documentation on zELDA are also available at https://zelda-ii.readthedocs.io/index.html. (redshift Estimator for Line profiles of Distant Lyman-Alpha emitters). To fit observed ${\rm Ly}\alpha$ line profiles, zELDA is now able to disentangle between the galactic and the IGM components of a ${\rm Ly}\alpha$ line profile through artificial neural network models trained with mock ${\rm Ly}\alpha$ spectra (see Gurung-Lopez et al. (2025) for a detailed description and testing of the algorithm or Sect. 2 for a summary).
This work aims to demonstrate on real data, i.e., actually observed spectra, how one can disentangle between the IGM and the ISM/CGM effects shaping the ${\rm Ly}\alpha$ spectral shapes. With this separation of effects at hand, we will also study the evolving impact of the IGM on spectral shapes, or vice versa study the IGM evolution purely using ${\rm Ly}\alpha$ spectra. A similar exercise has been done in previous work using stacked line profiles (Hayes et al., 2011). The novelty of this work is that this will be the first time doing so using individual spectra and estimating the ${\rm Ly}\alpha$ escape fraction in the separate sources. While the current work represents only the first step on utilizing this potential treasure trove of information (and thus, is to some degree a proof-of-concept study), other applications such as spatial tomography of the IGM are thinkable.
This work is structured as follows: in Sect. 2 we briefly introduce zELDA’s pipeline for recovering the galactic component from IGM attenuated ${\rm Ly}\alpha$ line profiles. We introduce the observational data set used in this work in Sect. 3. Next, we apply zELDA’s methodology to the observational data set in Sect. 4 and measure the mean IGM ${\rm Ly}\alpha$ escape fraction as a function of redshift. In Sect. 5 we validate our results by comparing them to the literature. Finally, we draw our conclusions in Sect. 6.
In this work, we draw the intergalactic transmission curves and ${\rm Ly}\alpha$ line profiles in the rest frame difference to the ${\rm Ly}\alpha$ wavelength ($\Delta\lambda_{0}$). Redshift accuracy is provided in the same units. Typically, this quantity is expressed in velocity units, instead, as $\Delta v=c\Delta\lambda_{0}/\lambda_{\rm Ly\alpha}\approx(247\,{\mathrm{km}\,\mathrm{s}^{-1}})\times\Delta\lambda/1$Å, where $c$ is the speed of light in vacuum and $\lambda_{\rm Ly\alpha}\approx 1215.67$Å.
Furthermore, we refer to the ${\rm Ly}\alpha$ IGM escape fraction of a source as the ratio between the flux emerging from the ISM/CGM and that observed after the IGM along the observer’s line of sight. This definition can be understood as the transmission of IGM around ${\rm Ly}\alpha$ and through the line of sight, convoluted with the line profile shape. In general, in the IGM, ${\rm Ly}\alpha$ photons are not destroyed by dust grains, although they are scattered out of the line of sight. This effectively causes a reduction of the observed ${\rm Ly}\alpha$ flux.
2 Reconstructing attenuated Lyman-$\alpha$ emission lines
In this section, we describe our approach to reconstructing observed ${\rm Ly}\alpha$ line profiles attenuated by the intergalactic medium (IGM) and to estimating their ${\rm Ly}\alpha$ IGM escape fraction. Our methodology is based on the Python package zELDA. Here, we briefly summarize zELDA’s pipeline, which is fully detailed in Gurung-Lopez et al. (2025), hereafter Z25.
To fit the observed ${\rm Ly}\alpha$ line profiles, we use the open-source Python package zELDA (Gurung-López et al., 2019a, 2021, 2022). zELDA ’s thin-shell model is implemented via linear interpolation within a precomputed grid of line profiles generated by the full Monte Carlo radiative transfer code LyaRT (Orsi et al., 2012). The grid contains 3,132,000 nodes spanning a five-dimensional parameter space: bulk expansion velocity $V_{\rm exp}$ $\in[0,1000]\,\mathrm{km\,s^{-1}}$, neutral hydrogen column density $N_{\rm H}$ $\in[10^{17},10^{21.5}]\,\mathrm{cm^{-2}}$, dust optical depth $\tau_{a}$ $\in[0.0001,1.0]$, intrinsic equivalent width $EW_{\rm in}$ $\in[0.1,1000]\,\AA$ , and line width $W_{\rm in}$ $\in[0.01,6]\,\AA$ of the ${\rm Ly}\alpha$ emission before entering the thin shell. Further details on the grid and its accuracy can be found in Gurung-López et al. (2022).
The IGM transmission curves used to train the IGM+z and IGM-z models are described in detail in Byrohl & Gronke (2020); Byrohl et al. (2021). These were computed within the TNG100 simulation (Naiman et al., 2018; Nelson et al., 2019; Marinacci et al., 2018; Pillepich et al., 2018; Springel et al., 2018). The ${\rm Ly}\alpha$ radiative transfer was calculated using an updated version of the ILTIS code (Behrens et al., 2019; Byrohl & Gronke, 2020; Byrohl et al., 2021). For each halo more massive than $5\times 10^{9}M_{\odot}$, transmission curves were obtained along 1000 random sightlines, at six redshift snapshots (z=0.0, 1.0, 2.0, 3.0, 4.0, 5.0z=0.0,\,1.0,\,2.0,\,3.0,\,4.0,\,5.0).
zELDA incorporates three models based on artificial neural networks (ANNs). Each model takes as input an observed ${\rm Ly}\alpha$ line profile. Both IGM+z and IGM-z account for IGM attenuation and aim to recover the unattenuated best-fitting shell model. IGM+z is trained to capture the redshift evolution of the IGM according to TNG100, while IGM-z is constructed to be redshift-independent. These two models therefore complement each other when measuring redshift-dependent properties. The third model, NoIGM, completely ignores IGM effects and fits a shell model directly to the observed line profile.
The three artificial neural network models are trained with mocked ${\rm Ly}\alpha$ line profiles produced with zELDA and the IGM transmission curves from Byrohl & Gronke (2020); Byrohl et al. (2021). The mock line profiles cover homogeneously the outflow parameters of the shell model. The training sets cover a wide range of spectral quality, from resolution of $W_{\rm g}$ =0.1Å to $W_{\rm g}$ =4.0Å and signal to noise of the peak of the line between $S/N_{p}$ =5.0 and $S/N_{p}$ =20.0. IGM+z, IGM-z, and NoIGM use a different training set and input. In summary:
  • IGM+z: The training set includes mock line profiles attenuated by the IGM. The IGM transmission curve associated with a particular ${\rm Ly}\alpha$ line profile uses the actual redshift of the source. By construction, the IGM evolution of Byrohl & Gronke (2020); Byrohl et al. (2021) is directly imprinted on the training set. Thus, ${\rm Ly}\alpha$ line profiles at lower redshift are less attenuated than those at high redshift. Finally, IGM+z includes a proxy redshift in the input.
  • IGM-z: The training set also includes mock line profiles attenuated by the IGM. Although, in contrast to IGM+z, we use the re-calibrated  [^2] As shown in Z25, Byrohl & Gronke (2020) IGM transmission curves were fine-tuned so that they match the mean evolution of the redshift of the IGM transmission ($\langle T_{\rm IGM}\rangle$) by Faucher-Giguère et al. (2008). Thus, the IGM transmission curves used in IGM+z and IGM-z are slightly different. transmission curves. These are drawn randomly without considering the mock source redshift. Thus, there is no evolution of the IGM through redshift in this training set. There will be ${\rm Ly}\alpha$ basically unattenuated (typical of $z=0$) and greatly attenuated (typical of high $z$) at all redshifts. Finally, unlike IGM+z, IGM-z does not include the proxy redshift in the input.
  • NoIGM: The training set does not use IGM transmission curves. NoIGM includes a proxy redshift in the input, as in IGM+z.
All three models provide the same type of output: the shell-model parameters with uncertainties, the estimated redshift, and the ${\rm Ly}\alpha$ IGM escape fraction within $\pm 2\,\AA$ of the line center, $f_{\rm esc}^{4\AA }$ (for IGM+z and IGM-z). By definition, $f_{\rm esc}^{4\AA }$ is obtained as the ratio of the IGM-convolved ${\rm Ly}\alpha$ flux to the intrinsic flux. Since the IGM at $z<4$ primarily affects the blue side of ${\rm Ly}\alpha$, $f_{\rm esc}^{4\AA }$ specifically measures absorption in the $\pm 2\,\AA$ window, and thus, while related, differs from the mean IGM transmission.
The use of both IGM+z and IGM-z enables us to test the robustness of our measurements. IGM+z encodes the expected redshift evolution of IGM transmission, whereas IGM-z does not. Comparing them allows us to identify potential biases in redshift-dependent trends, such as spurious variations in the IGM escape fraction induced by the training set. If IGM+z predicts a redshift trend absent in IGM-z, the effect is likely artificial and it could be attributed to our training set rather than being authentic. For reference, we also include results from NoIGM, which ignores IGM attenuation entirely.
This paper aims to evaluate how the ${\rm Ly}\alpha$ IGM escape fraction evolve with redshift. While the mean IGM transmission is well known to evolve, the redshift dependence of the ${\rm Ly}\alpha$ escape fraction has not been directly measured before. On the other hand, $\langle f_{\rm esc}^{4\AA }\rangle$ depends on the local IGM environment (typically denser and more neutral than the field) as well as on the ${\rm Ly}\alpha$ line profile shape. Thus, it is unclear whether the IGM+z training set reproduces its evolution accurately. We therefore include the IGM-z model as a redshift-independent control.
When the spectral quality is sufficiently high, both IGM-z and IGM+z recover $f_{\rm esc}^{4\AA }$ and the outflow parameters reliably, as shown in Z25. At low quality, however, the models behave differently due to their training sets. In particular, as quality decreases, the ANNs tend to return values that reflect prior distributions rather than true inference. Since IGM+z is forced to encode redshift evolution, it may reproduce trends even when absent in the data, while IGM-z will not. The close agreement between their predictions in the results section is reassuring, as it indicates that the information content of the observed spectra is sufficient for accurate inference rather than being dominated by training priors.
3 Description of observational data
The observational ${\rm Ly}\alpha$ line profiles used in this work are stored in the Lyman Alpha Spectral Database (LASD, Runnholm et al., 2021) [^3] http://lasd.lyman-alpha.com, which currently contains 349 ${\rm Ly}\alpha$ emission lines in the redshift range $0<z<6.0$. We exclude from our analysis:
  1. 1.
[…]
  2. 2.
    spectra with low $S/N_{p}$, and
  3. 3.
    spectra that apparently contain ${\rm Ly}\alpha$ emission from multiple sources.
After these cuts, our final sample comprises 313 ${\rm Ly}\alpha$ line profiles spanning $0<z<6.0$. All spectra are shown in Appendix F.
The ${\rm Ly}\alpha$ line profiles from LASD were obtained mainly from two experiments: i) the Cosmic Origins Spectrograph (COS Green et al., 2012) on board the Hubble Space Telescope (HST) and ii) the MUSE-WIDE survey (Urrutia et al., 2019; Herenz et al., 2017). ${\rm Ly}\alpha$ line profiles obtained via HST were observed in the General Observers (GO): GO 11522 and 12027 (PI: Green, Salzer et al., 2001; Wofford et al., 2013), GO11727 and 13017 (PI: Heckman, Heckman et al., 2011, 2015), GO 12269 (PI: Scarlata, Songaila et al., 2018), GO 12583 (PI: Hayes Hayes et al., 2014; Rivera-Thorsen et al., 2015), GO12928 (PI: Henry Henry et al., 2015), GO 13293 and 14080 (PI: Jaskot Jaskot & Oey, 2014; Jaskot et al., 2017), GO 14201 (PI: Malhotra Yang et al., 2017) and GO 13744 (PI: Thuan Izotov et al., 2016, 2018, 2020).
HST-COS and MUSE-WIDE exhibit very different spectral quality. The spectra obtained by HST have excellent spectral quality ($R\sim 16,000$). In particular, the ${\rm Ly}\alpha$ line profiles exhibit excellent wavelength sampling, as $\Delta\lambda_{\rm Pix}$ might take two particular values: 0.0598Å and 0.0735Å, depending on the medium-resolution gratings G130M and G160M. The spectral resolution is also of high quality. $W_{\rm g}$ and ranges from 0.073Å to 0.10Å with median 0.085Å. The signal-to-noise ratio of the maximum of the line, $S/N_{p}$, spans a wide range from $S/N_{p}$ =6.5 to $\sim 400$ with a median of $\sim$38 and only a $\sim 11\%$ of the sample exhibit $S/N_{p}$ <15.0<15.0. Considering MUSE data, the spectral quality is more limited ($R\sim 3000$), $\Delta\lambda_{\rm Pix}$ =1.25Å. Meanwhile, $W_{\rm g}$ changes from 1.25Å at $z=3$ to 4.0Å at $z=6.0$. We only consider ${\rm Ly}\alpha$ line profiles such as $S/N_{p}$ >>5.0. The maximum $S/N_{p}$ is 35.0, and the median is 7.8.
Figure 2: Nine examples of zELDA’s prediction on line profiles displayed in the rest frame. The observed line profile is shown in dark grey with its $1\sigma$ uncertainty in light grey. zELDA’s reconstruction using the models IGM+z, IGM-z, and NoIGM is displayed in blue, green, and yellow, respectively. The redshift of the source is shown in the top right corner. zELDA’s estimations of $f_{\rm esc}^{4\AA }$ given by the IGM+z and IGM-z models are shown in blue and green, respectively. The central grey-shaded region shows the wavelength interval of $f_{\rm esc}^{4\AA }$. The estimated outflow parameters are indicated in the table with each subplot. $V_{\rm exp}$ is given in  km s-1, $N_{\rm H}$ in $cm^{-2}$ and $EW_{\rm in}$ and $W_{\rm in}$ in Å. All the fitted line profiles are also shown in Appendix F.
In Z25, different spectral-quality configurations were explored. Most HST line profiles correspond to the best-case scenarios described there, while the MUSE spectra match intermediate and worst-case configurations. It was demonstrated in Z25 that zELDA can successfully reconstruct stacked ${\rm Ly}\alpha$ line profiles for both MUSE and HST qualities, and can reliably track the redshift evolution of $f_{\rm esc}^{4\AA }$.
The $S/N_{p}$ thresholds applied here were chosen to ensure accurate recovery of the stacked line profile, $f_{\rm esc}^{4\AA }$, and its mean $\langle f_{\rm esc}^{4\AA }\rangle$ across redshift, as shown in Z25. Specifically, for HST sources zELDA recovers the stacked line profile with a Kolmogorov–Smirnov (KS) estimator better than 0.04.444The KS estimator is computed between the reconstructed and true intrinsic ${\rm Ly}\alpha$ line profiles for 10,000 mock spectra spanning the full range of zELDA outflow parameters, IGM absorption levels, and spectral-quality configurations. For MUSE sources the stacked line profile is recovered with KS <0.11<0.11. As shown in Z25 (Sect. 4.3), with the spectral qualities of both MUSE and HST the redshift evolution (or lack thereof) of the stacked ${\rm Ly}\alpha$ profile is recovered robustly. Regarding $\langle f_{\rm esc}^{4\AA }\rangle$, for HST data both IGM+z and IGM-z retrieve unbiased values within 3%, whereas for MUSE-quality spectra $\langle f_{\rm esc}^{4\AA }\rangle$ may be biased by up to 10%. [^5] The precision of $f_{\rm esc}^{4\AA }$ estimates is evaluated using mock ${\rm Ly}\alpha$ profiles generated with zELDA under varying spectral qualities and IGM absorption levels.
[…]
4 Results on observed Lyman-$\alpha$ line profiles
In this section, we apply zELDA’s ANN models to observed ${\rm Ly}\alpha$ line profiles. In Sect. 4.1 we present a few examples of the predicted ${\rm Ly}\alpha$ line profiles. We study the correlation between $f_{\rm esc}^{4\AA }$ and outflow properties in Sect. 4.2. We present our observed $\langle f_{\rm esc}^{4\AA }\rangle$ reconstruction in Sect. 4.3.
4.1 Individual reconstructed ${\rm Ly}\alpha$ line profiles
In Fig. 2, we present nine examples illustrating the performance of zELDA on observed ${\rm Ly}\alpha$ line profiles. The observed spectra are shown in grey, while the predictions of IGM+z, IGM-z, and NoIGM are displayed in blue, green, and yellow, respectively. The top row corresponds to HST sources (A, B, and C), whereas the middle and bottom rows show MUSE LAEs (DI).
Overall, the three ANN models reproduce the red peak of the observed ${\rm Ly}\alpha$ line profiles well. Both IGM+z and IGM-z also predict the presence of a blue peak, while NoIGM tends to follow the observed spectra more closely without introducing additional features. This behavior is consistent across both HST and MUSE data. [^6] Note that in mock ${\rm Ly}\alpha$ line profiles, high-$z$ spectra lacking an observed blue peak are not always reconstructed by IGM+z and IGM-z with a blue peak.
At low redshift (top row of Fig. 2), the three models predict nearly identical intrinsic profiles for sources A and B, with close agreement between the predicted and observed spectra in both the red and blue peaks. This consistency reflects the negligible IGM attenuation at these redshifts. For source C ($z=0.36$), all models produce similar red peak shapes and amplitudes, but differ slightly in the blue peak: IGM-z predicts the strongest blue component, followed by IGM+z and then NoIGM. These differences correlate with the reconstructed expansion velocity, which decreases from 96.4 km s-1 (NoIGM) to 71.6 km s-1 (IGM+z) and 51.6 km s-1 (IGM-z). This trend suggests that including IGM attenuation impacts the inferred kinematic parameters, with IGM-z providing a more accurate fit by incorporating IGM constraints.
At higher redshift (middle and bottom rows), IGM absorption becomes more significant. For sources D, E, G, H, and I, both IGM+z and IGM-z predict a reconstructed blue peak—attenuated or partially absorbed—whereas NoIGM either predicts no blue peak or one with much lower amplitude. For instance, in source I, the suppressed blue peak predicted by IGM+z and IGM-z ($f_{\rm esc}^{4\AA }$ $\sim 0.56$) matches the observations, while NoIGM fails to reproduce this feature. This highlights the importance of IGM modeling for accurately reconstructing line profiles at high redshift. Moreover, the small difference in $f_{\rm esc}^{4\AA }$ values (within 5%) between IGM+z and IGM-z further underscores the robustness of their predictions.
The observed and reconstructed ${\rm Ly}\alpha$ line profiles of all the sources are shown in Appendix F. Fig. 19 for sources with $0.0<z<0.20$, Fig. 20 for sources with $0.20<z<0.35$, Fig. 21 for sources with $0.35<z<3.3$ , Fig. 22 for sources with $3.3<z<3.7$ , Fig. 23 for sources with $3.7<z<4.2$ ,Fig. 24 for sources with $4.2<z<4.8$ , Fig. 25 for sources with $4.8<z<6.0$. In general, the red peak of the ${\rm Ly}\alpha$ predicted by the ANN models follows the observations closely. The blue peak predictions follow the observations at low redshift, while at high redshift, they diverge for IGM+z and IGM-z.
Figure 4: $f_{\rm esc}^{4\AA }$ as a function of redshift for IGM+z (left) and IGM-z (middle). The right panel shows the histogram of the predicted $f_{\rm esc}^{4\AA }$ for IGM+z and IGM-z in blue an green, respectively.
4.2 Correlations between model and observed properties
[…]
We quantify the strength of the possible correlations by the Spearman correlation coefficients for the full sample, which are shown in Fig. 3, along with their $1\sigma$ uncertainties. Meanwhile, the full 1-dimensional and 2-dimensional distributions can be found in Appendix A. The coefficients from IGM-z are shown in the top-right corner (green, $\rho$), and those from IGM+z in the bottom-left corner (blue, $\rho_{z}$). Uncertainties are estimated following Gurung-López et al. (2022): we draw 1000 realizations of the outflow parameters from the ANN posteriors (see methodology in Z25), compute the Spearman correlation coefficient for each realization, and define the $1\sigma$ interval from the 16th and 84th percentiles of the resulting distribution. We consider two properties to be correlated (or anti-correlated) if the absolute value of the Spearman coefficient exceeds 0.45. Correlation matrices for the HST and MUSE subsamples separately are provided in Appendix D.
Overall, we find that the correlation coefficients inferred from IGM+z and IGM-z are very similar: both models typically identify the same correlations. In Appendix D, we present the coefficients for the low-redshift (HST, left) and high-redshift (MUSE, right) samples (see Fig. 17). Across the full dataset, only a handful of parameters show significant correlations, some of which are present in both subsamples (low and high redshift), while others appear only in one. These relations are the following:
[…]
Figure 5: Mean $f_{\rm esc}^{4\AA }$ as a function of redshift (squares) for IGM+z (blue) and IGM-z (green). The solid lines indicate the fits, including all the sources below $z=5$. Meanwhile, the colored dashed line indicates the fit using only MUSE data below $z=5$.
Table 1: Best-fitting parameters for the mean ${\rm Ly}\alpha$ IGM escape fraction ($\langle f_{\rm esc}^{4\AA }\rangle$) as a function of redshift, modeled with the modulated Fermi-Dirac distribution given by Equation 1.
$z$ range Experiments ANN model $a$ $b$ $c$
[…]
4.3 IGM escape fraction evolution with redshift
In this section, we examine in detail the redshift dependence of the IGM escape fraction, $f_{\rm esc}^{4\AA }$, as predicted by the models that include an IGM component (IGM+z and IGM-z).
As shown in Z25 (Sect. 4.3 and Appendix C), the accuracy of $f_{\rm esc}^{4\AA }$ depends strongly on the spectral quality of the ${\rm Ly}\alpha$ line profiles. We therefore restrict our analysis to spectra with $S/N_{p}$ >8.0>8.0 and $W_{\rm g}$ <3.0<3.0Å. This selection yields 93 HST and 113 MUSE sources, for a total of 206 ${\rm Ly}\alpha$ line profiles. These cuts ensure that the uncertainty in $f_{\rm esc}^{4\AA }$ remains below 0.05 for HST sources and below 0.09 for MUSE sources.
[…]
In Fig. 5, we present the mean ${\rm Ly}\alpha$ emission line IGM escape fraction, $\langle f_{\rm esc}^{4\AA }\rangle$, for IGM-z (blue squares) and IGM+z (green squares). Uncertainties correspond to the error of the mean, and the bins have been shifted by $\pm 0.01$ in redshift for clarity. At $z<0.5$ (HST data), IGM+z and IGM-z diverge slightly, with IGM+z predicting a constant value of $\sim 0.91$ and IGM-z $\sim 0.89$. At $z>3.0$ (MUSE data), however, both models agree closely, predicting that $\langle f_{\rm esc}^{4\AA }\rangle$ declines from $\sim 0.9$ at $z=2.9$ to $\sim 0.6$ at $z=5.0$.
In Fig. 5 we also show the best fitting of the predicted $\langle f_{\rm esc}^{4\AA }\rangle$ to a modulated Fermi-Dirac distribution (zELDA is able to recover the $f_{\rm esc}^{4\AA }$-$z$ evolution as discussed in Z25), i.e.,
$\langle f_{\rm esc}^{4\AA }\rangle=\displaystyle\frac{c}{e^{b(z-a)}+1}.$ (1)
The best-fitting parameters are listed in Tab. 1. For the full redshift range (HST+MUSE), the fits yield $c=0.93$ for IGM+z and $c=0.89$ for IGM-z. In the interval $2.9<z<5.0$, the parametric curves differ by less than 0.05. When restricted to MUSE data alone, both ANN models produce nearly identical fits, with $c=1.0$ and $a$ and $b$ consistent within $1\sigma$. This explains the differences at $z<2.0$ between the best-fitting curves derived from the full sample and from MUSE-only data. Extrapolating the fits, both models predict that the ${\rm Ly}\alpha$ IGM escape fraction approaches zero at $z\sim 11$.
[…]
5 Validation and discussion
In this section we aim to validate our results on the reconstructed ${\rm Ly}\alpha$ line profile shapes and $\langle f_{\rm esc}^{4\AA }\rangle$ redshift evolution. This task is made complex by an inherent aspect of Lya line profile reconstruction: we can only observe one sight line per source. In Z25 we calibrated and tested zELDA using multiple IGM transmission sight lines for the same ${\rm Ly}\alpha$ line profile, which is possible only for simulated spectra. We refer the reader to Z25 for an in-depth analysis of the performance of zELDA on mock ${\rm Ly}\alpha$ spectra with HST and MUSE quality. In order to show the robustness of the ${\rm Ly}\alpha$ line profile reconstruction and $\langle f_{\rm esc}^{4\AA }\rangle$ estimation we performed the following tests:
  • We show that the inferred redshift is highly accurate ($\sim$0.3Å) , with IGM-z being the least biased (see Sect. 5.1).
  • We show that the red tail of the ${\rm Ly}\alpha$ line profile, supposedly unaffected by the IGM at low redshift, is well recovered, specially when implementing the IGM absorption (see Sect. 5.2).
  • We computed the ${\rm Ly}\alpha$ stacked line profiles from the reconstructed ${\rm Ly}\alpha$ line profiles (see Sect. 5.3). We find that the stacked ${\rm Ly}\alpha$ line profiles do not evolve strongly with redshift. This indicates that the evolution in the observed ${\rm Ly}\alpha$ spectrum arises from the IGM absorption, in agreement with previous findings in the literature (Hayes et al., 2021a, 2023).
  • We demonstrate that our $\langle f_{\rm esc}^{4\AA }\rangle$ measurement in individual ${\rm Ly}\alpha$ line profiles is compatible with the IGM mean optical depth found in the literature, especially at $z>2$ (see Sect. 5.4) .
  • We find that our measurement of the IGM ${\rm Ly}\alpha$ escape fraction is consistent with the global ${\rm Ly}\alpha$ escape fraction reported in the literature in the whole redshift range (see Sect. 5.5).
Finally, we discuss the fact that the IGM+z and IGM-z models predict that the ${\rm Ly}\alpha$ line profiles are IGM attenuated at $z\sim 0$ in Sect. 5.6.
5.1 Accuracy of redshift estimates with zELDA
In this section, we evaluate the redshift accuracy of zELDA. For this purpose, we use HST sources with systemic redshifts measured from features other than ${\rm Ly}\alpha$, yielding a total of 111 sources. We focus on HST data because MUSE sources in LASD lack systemic redshift measurements below $z=0.5$. To quantify the precision, we analyze the distribution of rest-frame wavelength differences between the redshift predicted by zELDA ($z^{\rm zELDA}$) and the systemic redshift ($z^{\rm Sys}$), computed as in Gurung-López et al. (2022):
$\Delta\lambda^{\rm zELDA}_{0}=\lambda_{\rm Ly\alpha}\frac{z^{\rm zELDA}-z^{\rm Sys}}{1+z^{\rm Sys}},$ (2)
[…]
Our analysis shows that the three models IGM+z and IGM-z and NoIGM perform similarly. All of them recover accurately the ${\rm Ly}\alpha$ wavelength location in the line profiles with an uncertainty of $\sim 0.3$Å. At low redshift, where the IGM has only a minor impact on the ${\rm Ly}\alpha$ line profile, the performance gain of IGM+z and IGM-z over NoIGM is modest. However, at higher redshifts the effect of the IGM becomes more significant, and we expect the IGM-inclusive models to substantially outperform NoIGM.
A small systematic bias is present across all models, with zELDA generally overestimating the ${\rm Ly}\alpha$ rest frame by approximately 0.25Å (equivalent to $\sim$61 km s-1). This overestimation aligns with findings from previous studies, such as Orlitová et al. (2018) and Li et al. (2021), and points to inherent limitations of the ‘thin shell model.’ It is noticeable that IGM+z and NoIGM are biased evenly while IGM-z shows a better performance.
5.2 Line profile fitting accuracy
Figure 6: Rest frame difference between zELDA’s ${\rm Ly}\alpha$ wavelength and the estimated ${\rm Ly}\alpha$ wavelength using spectral features different from ${\rm Ly}\alpha$ in HST sources. zELDA’s predictions are shown blue, green, and yellow for IGM+z, IGM-z, and NoIGM, respectively. The mean and standard deviation of the distributions are shown for each model.
Figure 7: Left: Distribution of the Kolmogórov-Smirnov (KS) estimator in the red tail of HST ${\rm Ly}\alpha$ line profiles for IGM+z, IGM-z, and NoIGM in thick blue, medium green, and narrow yellow, respectively. The width of the distribution is shown for each distribution, computed as the difference between the 84th and 16th percentiles and the median.
Figure 8: Comparison between stacked ${\rm Ly}\alpha$ line profiles. The observed stacked ${\rm Ly}\alpha$ line profile is shown in grey, while zELDA’s prediction using IGM+z, IGM-z, and NoIGM are shown in blue, green, and yellow. The stacked ${\rm Ly}\alpha$ line profile is computed in 6 redshift bins. The relevant redshift interval.
[…]
In Fig. 7, we show the distribution of the KS estimator for the IGM+z (blue), IGM-z (green), and NoIGM (yellow) in the red tails of the HST sources. We find that IGM-z and IGM+z outperform NoIGM. In particular, the median of the KS distribution for the models is 0.07 for IGM+z, 0.07 for IGM-z, and 0.09 for NoIGM. Meanwhile, NoIGM exhibits a little bit more dispersion than IGM+z and IGM-z. We conclude that the IGM+z and IGM-z models predict the red tail of the observed ${\rm Ly}\alpha$ line profile with the same precision.
5.3 Stacked Lyman-alpha line profile
In this section, we analyze the redshift dependence of the observed ${\rm Ly}\alpha$ stacked line profile shape (a similar analysis was done by Hayes et al., 2021a, 2023; we focus here on the reconstruction of the pre-IGM spectra using zELDA). We limit our analysis to ${\rm Ly}\alpha$ line profiles with $S/N_{p}$ >>9.0. This ensures that more than $\sim$85% of the ${\rm Ly}\alpha$ line profiles of HST are recovered with KS<<0.1. Meanwhile, as $W_{\rm in}$ changes from 1.25Å to 4Å in MUSE data, $\sim$85% and $\sim$65% of the line profiles should be recovered with KS<<0.1, respectively (see appendix D in Z25). $S/N_{p}$ >>9.0 was chosen as a compromise between reconstruction quality and number counts. We verify that similar results are obtained across different signal-to-noise thresholds ($S/N_{p}$ >>6.0, 7.0, 8.0, 10.0), with higher thresholds producing noisier stacked line profiles, while the overall trends remain consistent.
We compute the stacked ${\rm Ly}\alpha$ line profile from a sample of observed ${\rm Ly}\alpha$ line profiles as follows. First, we apply zELDA’s ANN model (IGM+z, IGM-z or NoIGM) to the line profiles to obtain the reconstructed ${\rm Ly}\alpha$ line profiles and the redshift of the source. All the line profiles are shifted with respect to their rest frame in a fixed wavelength array. The individual ${\rm Ly}\alpha$ line profiles are normalized so that their maximum reaches unity. Then, the stacked line profile is computed as the median flux in each wavelength bin.
In Fig. 8, we present the ${\rm Ly}\alpha$ stacked line profiles across six redshift bins: $0.0<z<0.5$ (94 sources), $3.0<z<3.5$ (19 sources), and $3.5<z<4.0$ (17 sources) in the top row (left to right), and $4.0<z<4.5$ (10 sources), $4.5<z<5.0$ (13 sources), and $5.0<z<6.0$ (5 sources) in the bottom row (left to right). To ensure a fair comparison, the ${\rm Ly}\alpha$ profiles predicted by the ANN models are computed at the same resolution ($W_{\rm g}$) as the observations.
In general, we find that the observed stack spectrum (shown in grey in Fig. 8) changes with redshift. At $0.0<z<0.5$ we find a faint but detectable continuum. Meanwhile, at higher redshifts, no continuum is detected. At $0.0<z<0.5$, the ${\rm Ly}\alpha$ stacked line profile exhibits a clear double peak line profile, with a prominent red peak that extends $\sim 2\AA$ {} redder than ${\rm Ly}\alpha$. The blue peak is weaker and also extends $\sim 2\AA$ {} bluewards ${\rm Ly}\alpha$. The same structure of the double peak is also apparent at $3.0<z<3.5$. Thus, the ${\rm Ly}\alpha$ stacked line profiles for $0.0<z<0.5$ and $3.0<z<3.5$ exhibit only minor differences. At $3.5<z<4.0$, the blue peak becomes less pronounced and eventually disappears for $z>4.0$, while the red peak remains consistently present across all redshift bins. This is consistent with the findings by Hayes et al. (2021b), i.e., these results are compatible with the scenario in which the ${\rm Ly}\alpha$ line profile is attenuated by the IGM progressively at higher redshifts and the pre-IGM spectra show little evolution. The IGM absorption at these redshifts mainly affects the blue side of ${\rm Ly}\alpha$ emission. This causes the main evolution in the ${\rm Ly}\alpha$ stacked line profile to be at the blue peak.
Again, the observed stacked ${\rm Ly}\alpha$ line profile at redshift $0.0<z<0.5$ and that at $3.0<z<3.5$ exhibit a very similar blue peak. This is consistent with our measurements of the $\langle f_{\rm esc}^{4\AA }\rangle$ and the redshift invariance of the galactic stack line profile. We find that $\langle f_{\rm esc}^{4\AA }\rangle$ changes little from $z\sim 0.5$ to $z\sim 3.0$. Therefore, if the galactic stacked line profile is the same at $z\sim 0.5$ to $z\sim 3.0$, then the observed stack should be very similar at these redshifts. Reassuringly, this is what we find. This would imply that the IGM absorption in the ${\rm Ly}\alpha$ line profiles would not evolve strongly from $z\sim 0$ to $z\sim 3$. In contrast, the mean IGM transmission changes significantly in this redshift range. $\langle T_{\rm IGM}\rangle$ evolves from $\sim 1$ at $z=0$ to $\sim 0.75$ at $z=3$ (Faucher-Giguère et al., 2008). Thus, our results would indicate that the stacked ${\rm Ly}\alpha$ line profile is not very sensitive to the IGM evolution in this redshift window.
Regarding the ${\rm Ly}\alpha$ stacked line profiles predicted by the ANN models, we find that all three (IGM+z, IGM-z, and NoIGM) generally reproduce the observed profile redwards of the ${\rm Ly}\alpha$ wavelength. However, notable differences arise on the blue side. As expected, the NoIGM model closely follows the observed profile, showing a weak but noticeable blue peak at $z<4.0$ that nearly disappears at $z>4.0$.
In comparison, IGM+z and IGM-z behave similarly. Overall, the ${\rm Ly}\alpha$ stacked line profile predicted by both ANN models remains unchanged through redshift. Both ANN models predict that the red peak becomes slightly narrower with increasing $z$, as found in the observations. However, both IGM+z and IGM-z predict that the amplitude of the blue peak is relatively constant and that it becomes slightly narrower at higher redshift. This suggests that the physical mechanisms shaping the ${\rm Ly}\alpha$ line profile do not evolve strongly through cosmic time.
In summary, our reconstruction of the pre-IGM spectra using stacks strongly supports the picture suggested by Hayes et al. (2021b) in which the evolution of the observed ${\rm Ly}\alpha$ line profile shape is mainly due to the IGM. This further validates our pre-IGM line profile reconstruction. While they use a similar dataset (obtained from LASD; Runnholm et al., 2021), the method is very different from ours. While they construct a mean IGM transmission from empirical models, we made our analysis on a basis of source per source procedure. The strength of our methodology in comparison to theirs is that zELDA provides the reconstructed spectrum of individual sources affected by very different IGM line of sights.
5.4 Comparison between $\langle f_{\rm esc}^{4\AA }\rangle$ and the IGM mean transmission
Figure 9: Left: illustration of the Lya emission line IGM escape fraction computation from the stacked line profile. The solid colored lines show the mean IGM transmission of Byrohl & Gronke (2020) in increasing redshift from purple to red. These curves are recalibrated to match the mean IGM transmission of Faucher-Giguère et al. (2008). The dashed green and blue line show the reconstructed stack line profiles using zELDA models (IGM+z in blue and IGM-z in green). The shaded region marks the wavelength window there the Lya emission line IGM escape fraction is computed. Right: Mean Lya IGM escape fraction as a function of redshift (squares) for IGM+z (blue). The solid lines indicate the fits, including all the sources. Meanwhile, the colored dashed line indicates the fit using only MUSE data. The grey dashed line shows the Faucher-Giguère et al. (2008) mean IGM transmission. The yellow circles show the mean Lya IGM escape fraction using the Faucher-Giguère et al. (2008) recalibrated IGM transmission curves applied to the reconstructed stacked spectrum (i.e., convolving the solid lines with the dashed lines from the left panel).
The connection between the average IGM transmission, $\langle T_{\rm IGM}\rangle$, and the mean IGM ${\rm Ly}\alpha$ escape fraction, $\langle f_{\rm esc}^{4\AA }\rangle$, is initially unknown. $f_{\rm esc}^{4\AA }$ quantifies the fraction of ${\rm Ly}\alpha$ radiation escaping from LAEs over intermediate IGM scales (few cMpc), after being perturbed by RT in the ISM and CGM. $\langle T_{\rm IGM}\rangle$ is derived from the spectra of extragalactic sources within a wavelength window located several angstroms blueward of the ${\rm Ly}\alpha$ line, measuring the IGM hundreds of cMpc from the source. As a first approximation, the IGM primarily affects the blue side of the ${\rm Ly}\alpha$ line. Consequently, the measurement of $f_{\rm esc}^{4\AA }$ around the ${\rm Ly}\alpha$ wavelength naturally distinguishes $\langle T_{\rm IGM}\rangle$ from $\langle f_{\rm esc}^{4\AA }\rangle$. In addition, $f_{\rm esc}^{4\AA }$ is obtained by convolving the ${\rm Ly}\alpha$ line profile with the IGM transmission along a sightline. Therefore, the shape of the ${\rm Ly}\alpha$ line profile is crucial for assessing $\langle f_{\rm esc}^{4\AA }\rangle$, whereas $\langle T_{\rm IGM}\rangle$ remains independent of it. For instance, given a constant IGM transmission, a line profile with a stronger red peak than blue peak yields a higher $f_{\rm esc}^{4\AA }$, and vice versa.
To compare our measurement of $\langle f_{\rm esc}^{4\AA }\rangle$ with previous works, we compute a synthetic estimate of $\langle f_{\rm esc}^{4\AA }\rangle$ based on the redshift evolution of $\langle T_{\rm IGM}\rangle$ from Faucher-Giguère et al. (2008). Specifically, we calculate the mean ${\rm Ly}\alpha$ escape fraction, $\langle f_{\rm esc}^{4\AA }\rangle$, that corresponds to the mean IGM absorption reported by Faucher-Giguère et al. (2008), by convolving the mean IGM transmission curves around ${\rm Ly}\alpha$ with the average ${\rm Ly}\alpha$ line profile as a function of redshift. For this, we adopt the IGM transmission shape from Byrohl & Gronke (2020), recalibrated to match Faucher-Giguère et al. (2008). For the mean ${\rm Ly}\alpha$ line profile emerging from the ISM/CGM, we use the reconstructed stacked profile discussed in 5.3, which includes all sources from $z=0$ up to $z=5$. We repeated the analysis using stacked profiles in redshift bins and obtained consistent results, as the stacked spectrum varies little with redshift.
In the left panel of Fig. 9, we show the mean IGM transmission curves as a function of redshift (solid colored lines). The reconstructed stacked spectrum between $z=0$ and $z=5$, as modeled by IGM+z and IGM-z, is displayed in dashed blue and green lines, respectively. We compute the synthetic $\langle f_{\rm esc}^{4\AA }\rangle$ as the ratio between the convolution of the mean IGM transmission and the stacked line profile, and the unabsorbed stacked profile, within a window of $\pm 2\AA$ {} around ${\rm Ly}\alpha$ (grey region).
The synthetic $\langle f_{\rm esc}^{4\AA }\rangle$ derived from the stacked line profile and mean IGM transmission curves is shown as yellow dots in the right panel of Fig. 9. The synthetic $\langle f_{\rm esc}^{4\AA }\rangle$ decreases with redshift because the stacked line profile emerging from the ISM/CGM remains constant, while the blue peak is increasingly absorbed at higher $z$ as $\langle T_{\rm IGM}\rangle$ decreases (dashed grey curve). We find excellent agreement between the synthetic $\langle f_{\rm esc}^{4\AA }\rangle$ and our measurements from the ${\rm Ly}\alpha$ line profiles (blue squares), for both the IGM+z and IGM-z (not shown) models. At $z>2$, the synthetic and measured $\langle f_{\rm esc}^{4\AA }\rangle$ values are consistent within $1\sigma$, indicating good agreement with $\langle T_{\rm IGM}\rangle$. At $z<1$, however, we find that the measured $\langle f_{\rm esc}^{4\AA }\rangle$ is $\sim 8\%$ lower than the synthetic $\langle f_{\rm esc}^{4\AA }\rangle$ obtained using Faucher-Giguère et al. (2008).
Considering the parametric estimation of the redshift evolution of $\langle f_{\rm esc}^{4\AA }\rangle$, we find that when only MUSE data are used (dashed blue line), there is very good agreement with the synthetic $\langle f_{\rm esc}^{4\AA }\rangle$, showing consistency with the $\langle T_{\rm IGM}\rangle$ evolution from Faucher-Giguère et al. (2008). However, when including the HST dataset, the parametric estimate lies $\sim 8\%$ below the synthetic $\langle f_{\rm esc}^{4\AA }\rangle$ at $z<1.0$, while remaining consistent at $z>2$.
This result indicates that zELDA predicts ${\rm Ly}\alpha$ line profiles to be more absorbed than expected given the mean IGM optical depth at $z<1$, whereas at $z>1$ the predictions match the literature $\langle T_{\rm IGM}\rangle$ values. In Z25, we tested biases and accuracy in recovering $\langle f_{\rm esc}^{4\AA }\rangle$ as a function of redshift, showing that both IGM+z and IGM-z correctly recover $\langle f_{\rm esc}^{4\AA }\rangle$ at low redshift for a range of $\langle f_{\rm esc}^{4\AA }\rangle$ $(z)$ relations. Therefore, our findings may indicate either a limitation of the shell model at low $z$, or that ${\rm Ly}\alpha$ line profiles are indeed more attenuated than expected at $z<0.5$. In Sect. 5.6, we show several examples of $z<0.5$ sources that may exhibit IGM-attenuated ${\rm Ly}\alpha$ spectra, supporting this interpretation.
For comparison, we also show $\langle T_{\rm IGM}\rangle$, the mean IGM transmission from (Faucher-Giguère et al., 2008), as a dashed grey line. At $z<0.5$, the IGM-z and IGM+z predictions of $\langle f_{\rm esc}^{4\AA }\rangle$ fall $\sim 0.1$ below $\langle T_{\rm IGM}\rangle$. The difference between $f_{\rm esc}^{4\AA }$ and $\langle T_{\rm IGM}\rangle$ can be attributed to stronger absorption around the ${\rm Ly}\alpha$ line center (Laursen et al., 2011b; Gurung-López et al., 2022). Even if the IGM transmission is close to unity a few angstroms away from ${\rm Ly}\alpha$, stronger absorption in the range $2\AA {}<\Delta\lambda<0.5\AA {}$ can affect the blue peak of $z=0$ lines. At higher redshift ($z>2.0$), both the best-fit parametric curves for the ANN models and the predicted $\langle f_{\rm esc}^{4\AA }\rangle$ lie above $\langle T_{\rm IGM}\rangle$. This is because most of the ${\rm Ly}\alpha$ flux resides in the red peak of the ${\rm Ly}\alpha$ line profile, which remains largely unaffected by the IGM up to $z=5.0$. Thus, although $\langle T_{\rm IGM}\rangle$ can be as low as $\sim 10\%$ at $z=5.0$, a large fraction of ${\rm Ly}\alpha$ photons still escape through the IGM, as they are already redshifted relative to the ${\rm Ly}\alpha$ wavelength by the ISM/CGM. However, this trend could also reflect a selection effect: in this work, we measure $f_{\rm esc}^{4\AA }$ in MUSE galaxies with strong observed ${\rm Ly}\alpha$ emission rather than in the full galaxy population. Consequently, flux-limited samples may yield $\langle f_{\rm esc}^{4\AA }\rangle$ values higher than those of the overall star-forming galaxy population.
5.5 Comparison between the ${\rm Ly}\alpha$ global and IGM escape fractions
Figure 10: Comparison between the redshift evolution of the IGM-only and global ${\rm Ly}\alpha$ escape fractions. The filled colored squares show the IGM ${\rm Ly}\alpha$ escape fraction ($\langle f_{\rm esc}^{4\AA }\rangle$) found by IGM+z up to $z=5.0$, while those empty show the results at $z>5$. The colored solid line show the best fit to $\langle f_{\rm esc}^{4\AA }\rangle$ as presented previously. The dashed red line shows the global ${\rm Ly}\alpha$ escape fraction estimation by Hayes et al. (2011). The grey symbols show different studies from the literature measuring the global ${\rm Ly}\alpha$ escape fraction and their uncertainty.
Figure 11: Six examples of zELDA’s prediction on observed line profiles displayed in rest frame. The observed line profile is displayed in dark grey with its $1\sigma$ uncertainty in light grey. zELDA’s galactic line profiles of IGM+z, IGM-z and NoIGM are displayed in blue, green and yellow, respectively. The redshift of the source and the outflow properties for each model are shown in the corresponding color. $V_{\rm exp}$ is given in  km s-1, $N_{\rm H}$ in $cm^{-2}$ and $EW_{\rm in}$ and $W_{\rm in}$ in Å.The central grey shaded region shows the wavelength interval of $f_{\rm esc}^{4\AA }$. The red arrows mark potential IGM features.
[…]
In Fig. 10, we show the $f_{\rm esc}^{\rm Lya}$ redshift evolution obtained by Hayes et al. (2011) (dashed red line) along with the individual measurements from the other studies in grey as described in the legend. Other studies, such as Goovaerts et al. (2024) (based on Thai et al. (2023)), also obtained relations between the redshift and $f_{\rm esc}^{\rm Lya}$ , finding compatible results with Hayes et al. (2011). Overall, it is found in the literature that $f_{\rm esc}^{\rm Lya}$ increases with redshift, being earlier galaxies less attenuated in ${\rm Ly}\alpha$. Normally, this is attributed to the dust in the ISM of galaxies, which is less abundant at high redshift (Hayes et al., 2011; Gurung-López et al., 2019b; Goovaerts et al., 2024). In particular, the different experiments find that $f_{\rm esc}^{\rm Lya}<5\%$ at $z<3.5$. Also, at $5<z<6$, Ouchi et al. (2008); Cassata et al. (2011); de La Vieuville et al. (2019) found a $f_{\rm esc}^{\rm Lya}$ below 0.2 with relatively low uncertainty of $\sim 0.1$. Meanwhile, with larger uncertainties, Sobral et al. (2018) found a $f_{\rm esc}^{\rm Lya}\sim 0.3$ at $4<z<6$ and Goovaerts et al. (2024) found $f_{\rm esc}^{\rm Lya}\sim 0.8$ at $z=6$. Given the uncertainties, all these studies are mutually compatible.
[…]
Moreover, we find that the IGM contributes only marginally to the global ${\rm Ly}\alpha$ escape fraction up to $z\sim 3$, where $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.8$ while the total $f_{\rm esc}^{\rm Ly\alpha}$ remains below 0.1. At higher redshift, as the IGM becomes increasingly dense, its role in regulating the global escape fraction becomes more significant. In particular, at $z=5.0$ zELDA predicts $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.5$, while $f_{\rm esc}^{\rm Ly\alpha}\sim 0.2$. This implies that at this epoch the ISM/CGM absorbs about 40% of the ${\rm Ly}\alpha$ radiation. Thus, our results indicate that at redshift 5.0 the IGM is already the main absorber of the ${\rm Ly}\alpha$ emission line.
Our results therefore suggest that the IGM escape fraction decreases steadily with redshift, at least up to $z=5$. If this trend continues beyond $z=5$, the IGM would become more dominant, controlling the global ${\rm Ly}\alpha$ escape fraction, while the relative impact of the ISM/CGM would diminish.
5.6 IGM absorption in ${\rm Ly}\alpha$ line profiles at $z<0.5$
Interestingly, zELDA predicts that the ${\rm Ly}\alpha$ line profiles at $z<0.5$ are already attenuated by the IGM, while the mean IGM transmission is close to unity at $z=0$. In sections 2 and 4, we have discussed the IGM+z and IGM-z models. These two models have very different training sets. On the one hand, IGM+z makes use of the IGM transmission curves from Byrohl & Gronke (2020) without the recalibration [^7] We refer here to the recalibration to the IGM transmission curves from Byrohl & Gronke (2020) so that they match the mean IGM redshift evolution by Faucher-Giguère et al. (2008). See Z25. matching the redshift of the source. On the other hand, IGM-z uses the IGM transmission curves after recalibration to match the mean IGM transmission in Faucher-Giguère et al. (2008), and the IGM redshift is randomized. These differences cause the $f_{\rm esc}^{4\AA }$ distribution to evolve differently with redshift in the training sets (see section 3 in Z25). Thus, the $f_{\rm esc}^{4\AA }$ predictions are different between these models. For example, the IGM-z model gives an unbiased $f_{\rm esc}^{4\AA }$ measurement from $z=0$ to $z=5.0$ for mocks with different $\langle f_{\rm esc}^{4\AA }\rangle$ redshift evolutions with $S/N_{p}$ =15.0 and $W_{\rm in}$ =0.1Å. Meanwhile, for the same spectral quality, IGM-z overpredicts $\langle f_{\rm esc}^{4\AA }\rangle$ by a 5% when $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.8$ at $z<1.0$, while at $z>1.0$ $\langle f_{\rm esc}^{4\AA }\rangle$ is recovered correctly (see section 4.3 in Z25).
Despite the differences between IGM+z and IGM-z, both models predict that the ${\rm Ly}\alpha$ line profiles at $z<0.5$ are attenuated by the IGM. IGM+z predicts a mean attenuation of $\sim 0.92$. Also, IGM-z predicts $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.89$. Both models predict individual values ranging from $\sim$0.8 to 1.0. If these lines are IGM attenuated, given the high signal-to-noise and resolution of HST ${\rm Ly}\alpha$ line profiles, a few individual absorbed systems should be visible, similar to the ${\rm Ly}\alpha$ forest.
In Fig. 11, we show six examples of ${\rm Ly}\alpha$ line profiles that could be showing IGM features at $z<0.5$. Example A is a double peak ${\rm Ly}\alpha$ line profile that has a very good signal-to-noise in individual pixels. At $\Delta\lambda_{0}\sim-3\AA$ {}, in the tail of the blue peak, there is an apparent absorption. In contrast, IGM+z, IGM-z, and NoIGM predict that the tail of the blue peak decreases smoothly. This kind of feature is not present in the shell model. Thus, this might be an IGM absorption. Moreover, in example A, the red tail is quite well fitted by the three models. However, the red peak observed spectrum falls faster than the models. This could indicate that the IGM is absorbing the core of the line. This is also visible in examples C and D.
Also, in example B we find a plateau of flux $\sim 0$ at the core of the line (from $\Delta\lambda_{0}=\sim-1.5\AA$ {} to 0.3Å) [^8] Note that zELDA has problems reproducing these wide absorption troughs at line center since we restrict ourselves to $T=10^{4}\,$K gas. Higher (effective) temperatures lead to a wider absorption and thus can alleviate the need for central IGM absorption. This is not the case for absorption features off-line-center also marked in Fig. 11. . In comparison, the continuum of this source is detected around $0.05$. Additionally, this line shows a wide red peak that extends up to $\Delta\lambda_{0}\sim 4\AA$ {}. Symmetrically, the blue peak seems to extend down to $\Delta\lambda_{0}\sim-4\AA$ {}. As in case A, example B shows a decrease in flux in the blue peak at $\Delta\lambda_{0}\sim-3\AA$ {}. Neither of these two features is found in the shell models and might be compatible with IGM absorption.
Furthermore, in cases C, D and F, we find that the blue peak exhibits small drops in flux. However, considering the signal-to-noise of individual pixels, these drops are significant. The three models IGM+z, IGM-z, and NoIGM match well the red peak. The models and observations match the ascension of the blue peak ($\Delta\lambda_{0}$ between -0.5Å and 0.0Å) well – especially in cases C and D. Also, the predicted model is slightly above the observations, while they match again at the tail of the blue peak ($\Delta\lambda_{0}$ between -2.0Å and -1.8Å).
Finally, example E is a double peak ${\rm Ly}\alpha$ line profile with very good signal-to-noise in individual pixels. We find that IGM+z and NoIGM models predict almost the same intrinsic line profile, while IGM+z predicts a line with the same red peak but with a more substantial blue peak. Focusing on the blue peak, we find that in the observed ${\rm Ly}\alpha$ line profile, there is more flux in the tail of the blue peak than in the IGM+z and NoIGM models. However, the blue tail is very well fitted by the IGM-z model. An absorbing system in the IGM could also cause this.
In summary, across examples A–F, potential signatures of IGM absorption are consistently observed in the blue peak, including small flux drops and deviations from the smooth profiles predicted by our shell models. While more sophisticated intragalactic ${\rm Ly}\alpha$ radiative transfer models might provide somewhat better fits without IGM attenuation, note that generally such models have similar problems in carving out steep absorption features due to the typical diffusive frequency redistribution smoothing spectra. Additionally, features such as steeper declines in the red peaks and wider absorption plateaus at line center could further hint at the impact of IGM on ${\rm Ly}\alpha$ line profiles. Hence, these findings highlight the IGM’s potential role in shaping ${\rm Ly}\alpha$ profiles down to low redshifts, particularly in regions where intragalactic radiative transfer models struggle to fully replicate the observed spectra.
6 Summary and conclusions
The observed ${\rm Ly}\alpha$ line profile is the result of the combination of the radiative transfer processes taking place in the interstellar, circumgalactic, and intergalactic mediums. In this work, we have disentangled the IGM from the ISM/CGM components in 313 observed ${\rm Ly}\alpha$ line profiles by using the python package zELDA.
The latest version of zELDA includes different artificial neural network models to recover the galactic component from an IGM attenuated ${\rm Ly}\alpha$ line profile. These models are trained using mock ${\rm Ly}\alpha$ lines computed using a grid of precomputed ’shell model’ (Gurung-López et al., 2022) with a full Monte Carlo radiative transfer treatment (LyaRT, Orsi et al., 2012) and the ${\rm Ly}\alpha$ IGM transmission published by Byrohl & Gronke (2020) based on TNG100 simulation (Nelson et al., 2019).
We have applied zELDA to 313 observed ${\rm Ly}\alpha$ line profiles, extracted from LASD (Runnholm et al., 2021). Of them, 111 come from HST/COS observations, and 202 from MUSE observations. zELDA’s results on observed ${\rm Ly}\alpha$ line profile are:
  • In general, we find that ${\rm Ly}\alpha$ line profiles are little attenuated in HST data (z<<0.5). Meanwhile, in MUSE data ($z>3.0$), both, IGM+z and IGM-z, predict that the blue side of ${\rm Ly}\alpha$ is attenuated. The ISM ${\rm Ly}\alpha$ line profiles predicted by our models frequently exhibit a stronger blue peak than that observed. This trend becomes prominent at higher redshifts. This suggests that ${\rm Ly}\alpha$ line profile reconstruction becomes more important at higher redshifts. Therefore, galaxy kinematic studies using the ${\rm Ly}\alpha$ line profile at high redshift might benefit from using the IGM reconstructed line profile rather than the observed line profile directly.
  • We quantify the ${\rm Ly}\alpha$ line profile IGM escape fraction, $f_{\rm esc}^{4\AA }$, for the HST and MUSE samples. We find that the mean $f_{\rm esc}^{4\AA }$ at $z<0.5$ is different depending on the ANN model. While IGM+z predicts $\langle f_{\rm esc}^{4\AA }\rangle$ =0.92, IGM-z predicts $\langle f_{\rm esc}^{4\AA }\rangle$ =0.89. As discussed, IGM+z tends to be biased towards 1.0 in this redshift range, so the measurement provided by IGM-z should be more reliable. Moreover, both ANN models agree well when determining $\langle f_{\rm esc}^{4\AA }\rangle$ in MUSE data. IGM+z and IGM-z predict that $\langle f_{\rm esc}^{4\AA }\rangle$ evolve from 0.87 at $z=3.0$ to 0.55 at $z=5.0$. This measurement can be helpful to obtain a measured ’IGM-free’ ${\rm Ly}\alpha$ luminosity function. Interestingly, our models predict that $z<0.5$ ${\rm Ly}\alpha$ line profiles can exhibit IGM features such as those of the ${\rm Ly}\alpha$ forest. We have shown six potential examples of this.
  • The validation analyses confirm that zELDA provides robust reconstructions of ${\rm Ly}\alpha$ line profiles and reliable estimates of the redshift evolution of the IGM escape fraction. The method achieves sub-angstrom redshift accuracy ($\sim$0.3Å). The reconstructed profiles reproduce the red tail of the emission line with high fidelity and, when stacked, show little intrinsic evolution with redshift, indicating that the observed spectral variations are primarily driven by IGM absorption, in agreement with previous works (Hayes et al., 2021a).
  • We find that our estimate of the ${\rm Ly}\alpha$ IGM escape fraction is consistent with both the mean IGM optical depth and the global ${\rm Ly}\alpha$ escape fraction reported in the literature. Our results suggest that at $z\sim 5$ the IGM is the main absorber of the ${\rm Ly}\alpha$ emission line, attenuating $\sim 60\%$ of the total flux, while the combined contribution of the CGM and ISM accounts for the remaining $\sim 40\%$. If the observed trend of a decreasing $\langle f_{\rm esc}^{4\AA }\rangle$ with redshift extends to higher redshifts, the IGM would become even more dominant than the ISM/CGM in determining the observability of ${\rm Ly}\alpha$ emitters.
In times to come, zELDA could facilitate remarkable scientific explorations. For instance, with a substantial sample of LAEs and spectra of sufficient quality, zELDA may enable the 3D mapping of the IGM ${\rm Ly}\alpha$ escape fraction. This mapping could permit the analysis of cross-correlations between $f_{\rm esc}^{4\AA }$ and LAEs, potentially shedding light on whether the large-scale characteristics of the IGM are linked to ${\rm Ly}\alpha$ observability (Zheng et al., 2011).
Acknowledgements
The authors acknowledge the financial support from the MICIU with funding from the European Union NextGenerationEU and Generalitat Valenciana in the call Programa de Planes Complementarios de I+D+i (PRTR 2022) Project (VAL-JPAS), reference ASFAE/2022/025. This work is part of the research Project PID2023-149420NB-I00 funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU. This work is also supported by the project of excellence PROMETEO CIPROM/2023/21 of the Conselleria de Educación, Universidades y Empleo (Generalitat Valenciana). MG thanks the Max Planck Society for support through the Max Planck Research Group, and the European Union for support through ERC-2024-STG 101165038 (ReMMU). DS acknowledges the support by the Tsinghua Shui Mu Scholarship, funding of the National Key R&D Program of China (grant no. 2023YFA1605600), the science research grants from the China Manned Space Project with no. CMS-CSST2021-A05, and the Tsinghua University Initiative Scientific Research Program (no. 20223080023).
Appendix A Distribution of outflow parameters for IGM+z and IGM-z
Figure 12: Comparison of the ANN output and LASD variables in the observed ${\rm Ly}\alpha$ line profiles predicted by the same models. $EW_{\rm obs}$ and $L_{\rm Ly\alpha}$ were obtained from LASD. $V_{\rm exp}$ is given in  km s-1, $N_{\rm H}$ in $cm^{-2}$, $EW_{\rm in}$, $EW_{\rm ob}$ and $W_{\rm in}$ in Å and $L_{\rm Ly\alpha}$ in $erg/s$. The blue dots show the values predicted by IGM+z while the green dots show those of IGM-z. The grey dots show the properties that include measurements by LASD. The histograms on the diagonal show the distribution of the parameters in the corresponding color for each model. The strength of the correlations is quantified in Fig. 3
Fig. 12 shows the predicted parameters plotted against each other: $V_{\rm exp}$, $N_{\rm H}$, $\tau_{a}$, $EW_{\rm in}$, $W_{\rm in}$, $z$, $f_{\rm esc}^{4\AA }$, $EW_{\rm ob}$, and $L_{\rm Ly\alpha}$, arranged from left to right and top to bottom. The upper-right triangle (green and grey) corresponds to the properties predicted by IGM-z, while the lower-left triangle (blue and grey) corresponds to IGM+z. The diagonal panels show the one-dimensional distributions of each parameter. The strength of the (anti)correlations is quantified in Fig. 3 and discussed in Sect.4.2.
Appendix B Comparison of Outflow Properties across Models
Figure 13: Comparison of the ANN output variables in the observed ${\rm Ly}\alpha$ line profiles predicted by different models. The out given by IGM+z is given in the X-axis and IGM-z in the Y-axis. HST ($z<0.55$) sources are shown in blue, while MUSE ($z>2.9$) sources are shown in red. The black dashed line marks the one-to-one relation. The Spearman coefficient for the full sample is shown in black (top), for HST sources in blue (middle), and for MUSE in red (bottom). The red and blue solid lines are the best-fitting first-degree polynomials for the high and low redshift samples. In the top row $V_{\rm exp}$, $N_{\rm H}$ and $\tau_{a}$ are shown from left to right. In the bottom row $EW_{\rm in}$, $W_{\rm in}$ and $f_{\rm esc}^{4\AA }$ are shown from left to right. The mean uncertainty for the shown properties is displayed in the bottom right corner.
Figure 14: Same as for Fig. 13 but comparing IGM+z (X-axis) against NoIGM (Y-axis). NoIGM does not predict a value for $f_{\rm esc}^{4\AA }$, so we the comparison between NoIGM and IGM+z for $f_{\rm esc}^{4\AA }$ is not available.
In Fig. 13, we compare the outflow properties and $f_{\rm esc}^{4\AA }$ predicted by the IGM+z (X-axis) and those predicted by IGM-z (Y-axis). HST sources (low redshift) are shown in blue, while MUSE sources (high redshift) are shown in red. We find that, in general, IGM+z and IGM-z predict the same outflow properties and $f_{\rm esc}^{4\AA }$. The typical Spearman correlation coefficient for the whole sample (shown in the top left corner in black) is above 0.8. For the low and high redshift samples, the output variables also correlate. We find that both models behave almost the same for $V_{\rm exp}$, $N_{\rm H}$, $\tau_{a}$, and $EW_{\rm in}$ at high and low redshifts. Meanwhile, at high redshift, the $W_{\rm in}$ and $f_{\rm esc}^{4\AA }$ values predicted by IGM+z and IGM-z are close to the one-to-one relation. Also, in the low redshift sample, for $W_{\rm in}$, although the Spearman correlation coefficient is high (0.82), the $W_{\rm in}$ values predicted by IGM-z are $\sim 0.25\AA$ above the IGM+z predictions. Finally, for $f_{\rm esc}^{4\AA }$ the low redshift Spearman correlation coefficient is also high (0.71), but IGM-z predicts slightly lower values than IGM+z.
In Fig. 14, we show a comparison between the output variables of IGM+z (X-axis) and those predicted by NoIGM (Y-axis). In general, we find that there is a Spearman correlation coefficient higher than $\sim 0.7$ for the whole sample. The Spearman correlation coefficient is greater in the low redshift sample than in the high redshift sample. For example, for $V_{\rm exp}$, $\rho_{z<3}=0.72$ while $\rho_{z>3}=0.69$ and for $N_{\rm H}$ $\rho_{z<3}=0.83$ while $\rho_{z>3}=0.64$. The IGM+z model predicts higher values for $V_{\rm exp}$ at high redshift. Meanwhile, IGM+z predicts lower values of $W_{\rm in}$ than NoIGM at high redshift.
The results in Fig. 14 reveal systematic biases in shell model fits that do not account for IGM absorption (e.g. Gronke, 2017). These models overestimate expansion velocities by $\sim 0.30$ dex and neutral hydrogen column densities by $\sim 0.25$dex. This occurs because the IGM primarily absorbs the blue side of the line – hence, when models exclude the IGM, they compensate by increasing both the galactic absorption’s blueshift ($V_{\rm exp}$) and the strength ($N_{\rm H}$).
Appendix C Possible biases when fitting an IGM attenuated line with a model without IGM
Figure 15: Distribution of the difference between the input shell model properties and those recovered for the mock with HST-like quality. In the top panels $V_{\rm exp}$, $N_{\rm H}$ and $\tau_{a}$ are shown from left to right. In the bottom panels, $W_{\rm in}$, $EW_{\rm in}$, and $f_{\rm esc}^{4\AA }$ are shown from left to right. The prediction using IGM+z, IGM-z, and NoIGM is shown in blue, green, and yellow, respectively. The dashed horizontal lines indicate the distribution’s mean, matching the color used in the model. The black dashed line is set at 0 for reference. In top right corner a table summing up the mean ($\mu$) and standard deviation ($\sigma$) is displayed for the three models.
Figure 16: Same as Fig. 15, but for the mocks with the spectral quality of MUSE.
The shell model has been widely used to successfully reproduce observed ${\rm Ly}\alpha$ line shapes, providing valuable insights into the physical properties of ${\rm Ly}\alpha$-emitting systems. However, previous work has often neglected the impact of IGM absorption, which can systematically alter the observed line profiles and bias parameter estimates. In this section, we investigate the possible biases induced by fitting IGM-attenuated ${\rm Ly}\alpha$ line profiles with models that do not account for IGM effects.
In this section, we study the possible biases induced by fitting an IGM-attenuated line with a model without IGM. For this goal, we use NoIGM as the model without IGM absorption. Meanwhile, we also compare the results for the explored models with IGM, IGM+z, and IGM-z.
The bias in the shell model parameters in a given ${\rm Ly}\alpha$ line profile population might depend on the distribution of actual shell model parameters intrinsic to the sample. To make our analysis realistic, we build mock ${\rm Ly}\alpha$ line profile samples from the shell model parameters measured in the observed ${\rm Ly}\alpha$ line profiles in Sect. 4. In particular, we produce a single line profile with the same shell model parameters as each of the observed line profiles with matching spectral quality. Then, we convolve the spectrum with an IGM transmission curve at the redshift of the source and measure the $f_{\rm esc}^{4\AA }$. If the measured $f_{\rm esc}^{4\AA }$ is within 1% of that original measure, we keep the IGM transmission curve. Otherwise, we draw another one until the criteria are met. This ensures that the $f_{\rm esc}^{4\AA }$ distribution in the mock is very similar to that obtained from the observed ${\rm Ly}\alpha$ line profile. As a result of this procedure, the sample size, redshift distribution, and true shell model parameters in our mock match those of the observed sample.
In Fig. 15, we show the distribution of the difference between the measured shell model parameters and the true ones for the mock line profiles of HST. Meanwhile, in Fig. 16, we show those with MUSE spectral quality. In the top right corner, we display the mean and standard deviation of the distributions. Ideally, if the methodology and spectral quality were perfect, the output should match exactly the input, thus all the differences should be null, and the mean and standard deviation should be zero.
Focusing on the HST-like sample (Fig. 15), the IGM+z and IGM-z models exhibit biases lower than 0.06 dex, in general. Meanwhile, the NoIGM model is unbiased at the level of IGM+z and IGM-z in $W_{\rm in}$ and $EW_{\rm in}$. In contrast, the mean of the distributions of NoIGM is biased towards lower values of $V_{\rm exp}$ (0.1 dex) and higher values of $N_{\rm H}$ (0.2 dex) and $\tau_{a}$ (0.3 dex).
Regarding the MUSE-like sample (Fig. 16), the IGM+z and IGM-z models are generally less biased than NoIGM. In particular, IGM+z and IGM-z exhibit a bias smaller than 0.1 dex in $V_{\rm exp}$, $\tau_{a}$, $EW_{\rm in}$, and $f_{\rm esc}^{4\AA }$. We find that IGM+z and IGM-z are biased by $0.15$ dex in $W_{\rm in}$. Moreover, we find that NoIGM is more biased than IGM+z and IGM-z, particularly in $V_{\rm exp}$ and $W_{\rm in}$. Interestingly, we find that NoIGM (0.15 dex) is slightly less biased than IGM+z and IGM-z ($\sim$0.23 dex) when estimating $N_{\rm H}$.
Comparing the HST-like and MUSE-like samples, we find that the NoIGM model is generally more biased than IGM+z and IGM-z. Additionally, we find that the standard deviation of the distribution of differences changes with the spectral quality. In general, we find that the width of the distribution is large in the MUSE-like sample. This direct effect of the lower spectral quality makes the line profile reconstruction more difficult.
Appendix D Property correlation in MUSE and HST
Figure 17: Same as Fig. 3 but only for HST sources in the left panel and MUSE sources in the right panel. The uncertainty is not shown for simplicity but is similar to the previous figure’s.
In Fig.17 we show the Spearman correlation coefficients for the HST (left) and MUSE (right) samples. We find that in general, all the correlations found in the complete sample (see Fig.3) are also present in the individual samples, such as the correlation between $EW_{\rm in}$ and $EW_{\rm ob}$.
Other correlations existing in the combined sample are only present in one of them. For example, the anti-correlation between $f_{\rm esc}^{4\AA }$ and redshift is present in the MUSE and combined sample. However, no trend is found within the HST sources.
Appendix E $f_{\rm esc}^{4\AA }$ inference by an alternative model
Figure 18: Distributions of $f_{\rm esc}^{4\AA }$ predicted values for HST (left) and MUSE ${\rm Ly}\alpha$ line profiles (right). The distribution for REC+z, IGM+z, and IGM-z are shown in purple, blue, and green, respectively. The vertical dashed colored lines indicate the distribution median for the corresponding model.
[…]
We find the same trends in the performance of IGM+z, IGM-z, and REC+z in the observed spectrum as in the mock spectrum. The distribution of the predicted $f_{\rm esc}^{4\AA }$ values of the observed ${\rm Ly}\alpha$ line profiles presented in Sect. 4 for REC+z (purple), IGM+z (blue,) and IGM-z (green) are shown in Fig. 18. The left panel shows HST sources at $z<0.55$, while the right panel shows MUSE sources at $z>2.9$. For MUSE sources, the three models show very similar $f_{\rm esc}^{4\AA }$ distributions with $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.7$ (horizontal dashed lines). Not only is the distribution in $2.9<z<6.0$ the same, but also the redshift evolution, although it is not explicitly shown. Meanwhile, considering HST sources, the IGM+z and IGM-z exhibit very similar $f_{\rm esc}^{4\AA }$ distribution, with $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.89$ and $0.91$ respectively. Meanwhile, REC+z predicts higher $f_{\rm esc}^{4\AA }$ values, as the $f_{\rm esc}^{4\AA }$ distribution peaks at $\sim 1$ with small dispersion. Actually, the minimum $f_{\rm esc}^{4\AA }$ predicted by REC+z is $\sim 0.9$. Consequently, the $\langle f_{\rm esc}^{4\AA }\rangle$ for REC+z is $\sim 0.98$.
The training set in an artificial neural network determines the fidelity of its output. In particular, biases and trends inherent to the training set are imprinted on the output. We find that REC+z gives a $\langle f_{\rm esc}^{4\AA }\rangle$ of $\sim 1$ at $z<0.5$ in contrast to IGM+z and IGM-z that give $\langle f_{\rm esc}^{4\AA }\rangle$ $\sim 0.9$. Meanwhile, the three models give very similar predictions at high redshift. Therefore, the low $z$ discrepancy comes from the training set of REC+z. The low $f_{\rm esc}^{4\AA }$ variance at $z<1$ in REC+z causes its predictions to be biased towards $f_{\rm esc}^{4\AA }$ =1 at low $z$, rather than being a real prediction.
Appendix F All fitted lines
In this appendix, we show all the fitted lines for both MUSE and HST data. Line profiles are sorted by redshift. Fig. 19 shows sources with $0.0<z<0.20$, Fig. 20 for sources with $0.20<z<0.35$, Fig. 21 for sources with $0.35<z<3.3$ , Fig. 22 for sources with $3.3<z<3.7$ , Fig. 23 for sources with $3.7<z<4.2$ ,Fig. 24 for sources with $4.2<z<4.8$ and Fig. 25 shows sources with $4.8<z<6.0$. The observed ${\rm Ly}\alpha$ line profile is shown in grey, while the prediction of IGM-z, IGM+z, and NoIGM are shown in blue, green, and yellow, respectively. Furthermore, $f_{\rm esc}^{4\AA }$ is shown in the top left corner in the same color as the model used, alongside the source redshift. The grey shaded band marks the wavelength region where $f_{\rm esc}^{4\AA }$ is estimated.
In Fig. 26, we show the ${\rm Ly}\alpha$ line profiles from LASD that were not considered for the analysis in this work. They removed these line profiles because i) they looked noisy or ii) they exhibited a very steep continuum not modeled in zELDA, or iii) the fits of IGM-z, IGM+z, and NoIGM looked unsuccessful. From the original sample of 349 spectra, 35 (10%) were excluded.
In general, at low redshift (Fig. 19 and 20) IGM+z, IGM-z and NoIGM predict similar intrinsic ${\rm Ly}\alpha$ line profiles. In sources 7, 40, and 111 IGM+z and IGM-z predict almost the same ${\rm Ly}\alpha$ line profile, and NoIGM predicts a different one. In sources 1, 14, 30, 38, 66, 82, and 97 IGM+z and IGM-z actually predict a different intrinsic ${\rm Ly}\alpha$ line profile. Also, in general, the $f_{\rm esc}^{4\AA }$ predicted by IGM+z and IGM-z are close to unity, with IGM-z $f_{\rm esc}^{4\AA }$ predictions slightly lower than those of IGM+z.
Meanwhile, at higher redshift (Fig. 22, 23, 24, and 25), the intrinsic ${\rm Ly}\alpha$ line profile predicted by IGM+z and IGM-z match, while they diverge from that of NoIGM. Although in some cases, the three models match relatively well, such as in sources 116, 119, 142, 162, 186, 220, and 299. Note that this scenario is more frequent at lower redshift, closer to $z=3.0$. Meanwhile, at higher redshifts, closer to $z=6.0$, there is more difference between the prediction and IGM+z/IGM-z and NoIGM. Also, in $\sim$80% of the high redshift sources, the IGM+z and IGM-z predict the presence of a blue peak in the ${\rm Ly}\alpha$ intrinsic spectrum.
Figure 19: zELDA’s prediction on observed line profiles displayed in restframe. The observed line profile is shown in grey. zELDA’s reconstruction using the models IGM+z, IGM-z, and NoIGM is displayed in blue, green, and yellow, respectively. The redshift of the source is shown in the top left corner. zELDA’s estimations of $f_{\rm esc}^{4\AA }$ given by the IGM+z and IGM-z models are shown in blue and green, respectively. The grey-shaded region shows the wavelength interval of $f_{\rm esc}^{4\AA }$. The red number indicates the index in LASD.
Figure 20: Same as Fig. 19, but for other observed ${\rm Ly}\alpha$ line profiles between $z=0.2$ and $z=0.35$.
Figure 21: Same as Fig. 19, but for other observed ${\rm Ly}\alpha$ line profiles between $z=0.35$ and $z=3.3$.
Figure 22: Same as Fig. 19, but for other observed ${\rm Ly}\alpha$ line profiles between $z=3.3$ and $z=3.7$.
Figure 23: Same as Fig. 19, but for other observed ${\rm Ly}\alpha$ line profiles between $z=3.7$ and $z=4.20$.
Figure 24: Same as Fig. 19, but for other observed ${\rm Ly}\alpha$ line profiles between $z=4.3$ and $z=4.8$.
Figure 25: Same as Fig. 19, but for other observed ${\rm Ly}\alpha$ line profiles between $z=4.8$ and $z=6.0$.
Figure 26: Same as Fig. 19, but for other observed ${\rm Ly}\alpha$ line profiles that exhibit low $S/N_{p}$ (e.g., 0 , 1 ), a steep continuum not modeled in zELDA (e.g., 0 , 1 ), or with more than two components (e.g., 3).